### v1

# CAMeleon: Interactively Exploring Craft Workflows in CAD

Abstract.
Designers of physical objects make assumptions on the material and fabrication workflow early in the design process. Recovering from bad assumptions is hard, because the design and resulting CAD model are locked-in to those assumptions. We present CAMeleon, a software tool to interactively explore fabrication workflows at any stage of the CAD process.
CAMeleon’s modular architecture allows users to execute their design with different workflows, and preview results. Users can freely explore alternative workflows. CAMeleon’s architecture, can be extended with new workflows, increasing the richness of workflows available.
We implemented five fabrication workflows in CAMeleon. We demonstrate CAMeleon’s extensibility through collaboration with six craftsmen whose workflows we replicated. We also implemented workflows of three papers and reflect on the process of these nine extensions. A usability study (n=12) showed that CAMeleon allowed participants to explore and select workflows for their design which they did not know before.
1. INTRODUCTION
Advances in computer-aided design (CAD) software, and computational fabrication techniques (e.g., additive manufacturing, subtractive fabrication, and wire forming) have enabled an increasingly broad set of people to engage in the design of physical objects. This trend towards personal fabrication (Baudisch et al., 2017) enabled marginalized communities to create custom assistive technology (Hurst and Tobias, 2011; Parry-Hill et al., 2017; Higgins et al., 2022), allowed high-schools to make fabrication part of their curriculum (Baudisch et al., 2019; Roumen et al., 2022; Hartikainen et al., 2024), and empowered several exciting hardware start-ups (e.g., Prusa3D, bambu, and XTOOL). As this trend continues, some envision fabrication machines to become as ubiquitous as kitchen appliances (Baudisch et al., 2023).
Designing physical objects presents challenges, particularly for people who are not well-versed in fabrication processes. Before even starting to design objects, designers are expected to make assumptions about the final fabrication process, as it has implications for how to design the structure supporting that process. While designers can bypass this challenge by obtaining shape models through 3D scanning, downloading, or through generative AI. The resulting models typically lack critical information about the fabrication processes needed to bring the designs to life, and changing the fabrication process to better fit the needs of the designer, is all but impossible.
To address these challenges, HCI researchers have developed tools to make users really good at designing with a specific process/material in mind. Tools like Kyub (Baudisch et al., 2019) and FlatFitFab (McCrae et al., 2014a) are specialized for laser cutting, Bridging the gap (Dumas et al., 2014) and Stress relief (Stava et al., 2012) for 3D printing, and Plushie (Mori and Igarashi, 2007) for textile design. These tools have made computational fabrication accessible by reducing the required upfront domain knowledge of fabrication processes and materials. The price of using such software is that designers lock themselves into one specific fabrication process from the get-go (e.g., designing for laser cutting or for 3D printing). Ironically, as designers make progress on their design, they learn more about limitations/constraints imposed by the fabrication processes they chose, while at the same time it gets ever harder to explore alternatives.
On the other end of the spectrum, there is the industrial workflow where engineers first design in CAD and then use so called ”Computer Aided Manufacturing” (CAM) to convert their design to machine-specific instructions. While CAM generates instructions for different machines, it still requires a design that is optimized for that specific fabrication process. These CAM workflows are heavily machine-oriented (e.g., they convert STLs to g-code for printing), and fail to acknowledge that most fabrication processes (especially craft workflows) involve manual processes.
In this paper, we propose CAMeleon, a software tool to interactively explore craft workflows in Computer Aided Design.CAMeleon, allows users to defer the decision of what fabrication workflow and material to use, and lets users interactively explore alternatives until they start fabricating. It achieves this by encapsulating craft and fabrication workflows as modular components of the interface, each taking volumetric shape as input, converting that to workflow-specific geometry, and exporting instructions for both machines and humans. We call these instructions human-centered CAM workflows. As the target volumetric shape remains intact, users can freely explore alternative workflows before making final decisions. The modular architecture of CAMeleon makes it easily extensible, further increasing the richness of workflows available to end-users.
We identify three core use cases for CAMeleon: (1) during design: supporting designers to explore and preview different workflows for their product, (2) after design or using generative AI: serving as a tool after design, in a generative AI process to create shape models, or when users browse shape models, and then explore different manifestations thereof through CAMeleon. (3) after fabrication: re-fabricating existing objects with different fabrication workflows through 3D scanning.
We surveyed common craft workflows shared online, we classified these into categories, and from each of these categories we implemented one workflow in CAMeleon. To study the extensibility of CAMeleon, in collaboration with six craft experts we replicated their workflows in CAMeleon, we furthermore implemented the workflows of three research papers. We reflect on the process of extending CAMeleon through implementation and through follow-up interviews with the expert craft experts. In a usability study with 12 participants, we found that participants were more likely to explore different fabrication processes, allowing them to learn about fabrication and craft.
We conclude that deferring the decision of fabrication processes, and making this decision interactive, allows designers to develop a more explorative attitude on workflows for computational fabrication. Especially designers with expertise that deviates from fabrication processes benefit from the ability to revise their decisions later on. As shape-based CAD models become more accessible, and tools like CAMeleon provide methods to capture human-centered CAM workflows, similar to recipes for cooking appliances, we reduce the barrier to entry for many more users to start taking advantage of their ability to create a physical reality that fits their needs.
2. WALK-THROUGH
Figure 2 illustrates the process of Peter creating a small stool for his child using CAMeleon. (a) Peter begins by downloading a 3D volumetric model (here an STL file) from Thingiverse, an online repository of 3D models. He uploads this model into CAMeleon. (b) Unsure of which workflow best meets his needs, Peter starts by exploring the available options. He browses through some options, and becomes familiar with the various workflows and the machines required for fabrication, displayed as icons beneath each workflow. Peter has access to a laser cutter, a 3D printer, and a wire bender, which allows him to consider different approaches for creating the stool. Given the need for a load-bearing structure with a smooth finish for his child’s safety, Peter explores and compares workflows that meet his requirements, he drags the ’Mold making,’ ’Wire mesh sculpture,’ ’Thermal Wire Cutting for Foam Fabrication’, and ’Epoxy coating’ workflows to the view-port. After comparing his options, Peter decides that the ’Epoxy coating’ workflow is the most suitable. (c) He changed the layer height to 3 mm to fit the material he has in store. Once he is satisfied with his design, Peter exports the machine-instruction files (here SVG for laser cutting).
After exporting, CAMeleon assists Peter step by step in executing the manual steps to build this stool. Figure 3 illustrates the interactive human-centered CAM workflow. (a) Peter already did the first step: laser cutting. (b) Peter then assembles and glues the cut pieces, fitting each layer together to form the structure. (c) Step 3: Peter completes the assembly by sanding the laser-cut structure to achieve a smooth finish. (d) Finally, Peter applies a layer of epoxy resin, carefully coating the entire structure for a sealed and polished finish.
3. CONTRIBUTIONS, BENEFITS, AND LIMITATIONS
We make three core contributions:
(1) The development of an interactive CAD tool, CAMeleon, that allows users to explore and compare fabrication workflows. We discuss the implementation and design decisions that went into the development of CAMeleon and release the code as an open source project for others to extend.
(2) The concept of human centered CAM workflows that include manual operations to capture realistic fabrication processes/workflows beyond mere machine instructions.
(3) The sample implementation of 14 fabrication workflows, including those derived from collaborations with craftsmen and recent research papers, to provide interactive fabrication guidance. We reflect on this extension process and what it takes for others to extend this further. This furthermore serves as a bridge for experts to distribute their workflows and processes across more users.
Limitations include that not every process can simply be converted into another, occasionally there are size or other technical requirements specific to a process that prevent conversion. Processes in CAMeleon include warnings to users for such requirements (e.g., ”features in this model are too small for PVC pipe bending, some details will get lost”) and its interactive preview gives them a sense of the expected outcome. Another limitation is that in the current state, it takes quite a bit of knowledge of CAMeleon’s code base and Three.JS to implement new workflows, in the future we plan to design a specific extension interface for CAMeleon to make this more accessible to craftsmen and people who invent new fabrication processes.
4. RELATED WORK
This work builds on research in software for personal fabrication, design for volumetric fabrication, and documenting workflows.
4.1. Software for Personal Fabrication
Modern fabrication processes, such as 3D printing, laser cutting, and CNC milling, have empowered users to personally fabricate custom objects (Baudisch et al., 2017). Software tools to support such fabrication processes have lowered the barrier to entry for novice users to create high-quality products. They typically support users with domain knowledge in a specific fabrication process. CAD tools for fabrication such as kyub (Baudisch et al., 2019), FlatFab (McCrae et al., 2014b) are specifically designed for one fabrication process, here laser cutting. Other tools like CoilCAM (Bourgault et al., 2023) focus on clay printing, or Imprimer (Tran O’Leary et al., 2023) for CNC milling take a more creative approach to allow users to experiment with these processes. They typically use a form of computational notebooks, such as p5.fab(Subbaraman and Peek, 2022) which allows exploration of processes across different techniques.
4.2. Design for Volumetric Fabrication
Fabricating volumetric shape often relies on various support structures, such as molds and jigs. Each serve different purposes such as facilitating shaping (e.g., plaster molds in ceramics, templates in woodworking), reinforcing structures (e.g., rebar molds in construction, wooden frameworks in shipbuilding).
Tools like WrapIt facilitate manual wire bending for jewelry by generating jigs from 2D patterns, enabling users to easily hand-bend wires into desired shapes (Iarussi et al., 2015). Fabricable 3D Wire Art generates jigs based on 3D models and images, enabling more complex designs (Tojo et al., 2024). In laser cutting, StackMold aids designers by generating molds for prototyping multi-material and multi-colored objects embedded with electronics, streamlining the fabrication process (Valkeneers et al., 2019). Schwartzburg et al.’s Intersecting Planar Pieces provides techniques for creating structures through precise cuts and component assembly, enhancing the versatility of 2D subtractive fabrication (Schwartzburg and Pauly, 2013). In the field of 3D printing, Schmidt et al. developed automated support generation systems that stabilize complex prints during 3d printing (Schmidt and Umetani, 2014). Innovations like CoreCavity offer interactive shell decomposition methods to generate two-piece rigid molds for diverse materials (Nakashima et al., 2018), while Metamolds produce 3D-printed containers designed for creating silicone molds, expanding the material possibilities for designers (Alderighi et al., 2018). Diverging from traditional support methods, Scrappy introduces a novel approach by utilizing scrap materials as infill for 3D prints, promoting sustainability and resource efficiency in fabrication processes (Wall et al., 2021).
These tools effectively assist in shaping during craft processes or converting existing models into complex structures, thereby lowering the technical barriers and making these processes more accessible. Building on this background, our goal is to create a tool whih allows users to explore and compare such workflows, allowing more people to access different techniques. This platform aims to bridge the gap between novice users and craftspeople, serving as both a resource for learning about craft techniques and an educational tool for sharing craft knowledge.
4.3. Documenting Workflows
In disciplines such as industrial manufacturing (Derhamy et al., 2018), and business process management (Goo, 2010), there has been a long standing focus on documenting workflows and processes. Such documentation has been instrumental to increase efficiency (Rai et al., 2009), detect and minimize errors (Schiff and Bates, 2010), and support continuous process improvement (Tibeme et al., 2018). It furthermore facilitates the transfer of knowledge and ideas across teams or enterprises (Zhao et al., 2018). In craft communities such workflow knowledge often is transferred from one to another and fails to benefit from similar distribution.
Some research has begun to collect and integrate multiple fabrication workflows into a single platform, making significant strides in advancing digital fabrication. For example, Vespidae offers flexibility in controlling fabrication machines, allowing them to manipulate key parameters. It features modular components that users can freely combine to form new code definitions, facilitating a deeper exploration of the design space (Fossdal et al., 2023). Similarly, Tandem addresses the complexity of experimental digital fabrication workflows, successfully capturing and recording workflows across multiple software tools, including machine tuning and manual interventions (Tran O’Leary et al., 2024). Research such as Process Products captures iterative processes by embedding this information directly into objects, which helps designers communicate design evolution and aids manufacturers in understanding tool usage (Tseng and Tsai, 2015). Similarly, Product versus Process explores how design documentation is created and used by authors and readers on platforms like Instructables111https://www.instructables.com/ (Tseng and Resnick, 2014).
On the hobbyist side of the spectrum, online communities such as Instructables, Dorkbot222https://dorkbot.org/ and ravelry333https://www.ravelry.com/, serve as knowledge-sharing hubs where makers showcase their projects. These platforms allow users to share detailed documentation of fabrication processes, including the use of machines, software, materials, and techniques, while engaging in discussions with other makers to exchange tips and experiences. On the other hand, most programming tools for fabrication primarily focus on the digital fabrication part of the process, without delving into how users can approach subsequent craft techniques or enhance their projects through hands-on methods. P5.fab (Subbaraman and Peek, 2022) make this case by opening up CAM workflows to more creative manipulation. For example, tools like WrapIt offer strong support in generating jigs for wire bending (Iarussi et al., 2015), EscapeLoom enhances the design space for weaving by building support structures (Deshpande et al., 2021), and Metamolds automatically defines the cuts needed to extract cast objects from silicone molds (Alderighi et al., 2018). While these tools excel in their respective areas, they could be complemented with resources that provide users guidance on the manual techniques required in these crafts.
5. Surveying craft workflows
The goal of CAMeleon is to explore craft workflows. To understand this space, we explored processes shared in online communities. We reviewed 150 winning projects from the ’Workshop’ category on Instructables444https://www.instructables.com (sorted by ’I Made Its’). We excluded workflows centered on microprocessor-based products, purely 3D printed items without any additional workflow, and those focused solely on polishing, to maintain an emphasis on fabrication workflows. This left us with 56 projects. The appendix contains a list of all workflows used for this process with links to their instruction pages. We proceeded to cluster these into categories to better understand the design space. In Figure 4, the different processes are shown and divided into the following categories: Wire Forming, Interlocking Structure, Stacked Slice Construction, Guide Structure, Mold Casting. Two workflows did not fit this categorization, one of them was creating shapes with epoxy resin and the other building boats from wooden spars and frames.
To start off with a representative subset of fabrication processes based on this classification, we picked one representative sample of each category as what we call foundational processes. Figure 5 shows an overview of these: (a) Thermal Wire Cutting for Foam Fabrication, (b) Mold Casting, (c) Laser-Cut Interlocking Structures, (d) Wire Mesh Fabrication, and (e) Stacked Slice Construction. The variety of these processes not only gives users five interesting workflows to start out with, but they also serve as good templates as building blocks for others to produce workflows. By building human-centred CAM workflows for these processes, we allow users to convert their design to be fabricated with these workflows and generate both machine instructions and manual guidance.
6. CAMeleon
In this section we discuss design considerations and technical details that form the basis of CAMeleon. We discuss the interface design, the implementation, and demonstrate how we designed the system to support three different use cases.
6.1. Interface Design
CAMeleon runs in user’s browser to maximize support across systems and architectures. The core of the interface is the way users explore different fabrication workflows, this is fundamentally supported by the comparison tool and the ability to browse different workflows. The other core aspect of the interface are the interactive guides to create human-centred CAM workflows. We discuss each individually in this subsection.
6.1.1. Comparison Windows
The CAMeleon comparison tool is designed to compare different workflows. The comparison tool lets users drag up to four workflows into the view-port, drawing inspiration from visual studio code editors. All workflows load the same 3D model, enabling users to clearly see and compare the potential outcomes side by side. This visual comparison helps users better understand how different workflows will influence the final product, making it easier to select the most suitable option based on their needs. Each window can be individually closed using the (x) icon on the top-right to free up space for other workflows or use more screen estate for the chosen workflow.
6.1.2. Workflow Browser & Filtering
To be able to compare different workflows, CAMeleon has a workflow browser panel on the left of the view-port. This is where users select workflows to be used in their fabrication process. Figure 7 shows different ways of using the browser. The most straight-forward is to scroll through the processes, however this obviously does not scale as the number of processes increases. To help them search more effectively CAMeleon lets them filter by keywords or by practical process-related dimensions.
Before categorizing the workflows in our survey, we placed them on different dimensions to compare and organize. These dimensions(e.g., what materials they use, what types of structure, or the degree of load-bearing capacity) formed an additional filter mechanism to find similar human centered CAM workflows. We found three categories of dimensions which are presented to users as shown in Figure 7(b).
Product-related Dimensions: These dimensions encompass factors such as load-bearing capacity, high-temperature tolerance, the need for a lightweight structure, and the level of fidelity required to reproduce intricate details. The choice of different support structures or materials can introduce unique features to the final product, enabling users to tailor their projects to specific needs.
Structure-related Dimensions: These dimensions focus on the use and construction of support structures, jigs, or molds. They address aspects like the removability of the support structure, its integration within the object, flexibility, modularity, and reusability. This categorization helps users understand and manage the specific requirements and characteristics of the support structures used in their fabrication process.
Machine-related Dimensions: The choice of fabrication machinery is closely linked to the selection of manufacturing processes. These dimensions guide users in identifying and selecting workflows based on the machines required, such as 3D printers, laser cutters, or wire benders.
6.1.3. Interactive Guide for human-centred CAM workflows
As mentioned before, traditional CAM tools focus on machine-instructions, and fail to deliver instructions to humans who use the resulting artifacts and/or control the machines. We therefore propose human-centred CAM workflows which contains machine-instructions complemented with interactive guidance for users.
Figure 8 shows such guidance, click through the different steps which allows them to follow the assembly process at their own pace and go back if they missed a step. (a) This assembly guide includes textual information, detailing how to assemble each component, with tracks and tips highlighted to ensure clarity. (b) Additionally, the guide includes links to external resources, such as YouTube tutorials, for in-depth guidance on the specific technique or examples of similar processes. This interface is designed to be similar to Instructables pages to slot in with existing experience of craftspeople and makers.
6.2. Implementation
CAMeleon is an open-source tool using Node.js and Three.js for 3D visualization and mesh processing. Three.js is used for its efficient handling of 3D models in the browser, while Node.js ensures scalable back-end support for file management and workflow coordination.
The tool processes STL files (containing triangle meshes), modifying the mesh and generating outputs like STL, SVG, or CSV depending on what machine is used. The system follows a modular architecture, where each workflow is independently defined in its own HTML file, allowing new workflows to be added without altering the core structure. The index.js file acts as a controller, dynamically loading and executing workflows based on user input. Each workflow follows a standardized pattern: receiving an STL file, processing it through Three.js libraries, and exporting the desired format.
6.3. Use Cases
We designed CAMeleon to work for three different categories of usage: during the design process, after the design process, and after fabrication.
During Design. During an ongoing design process, in CAD tools like Fusion360, users may want to explore how their design decisions impact the final fabrication workflow, or how different workflows can guide them in their design. To support this, we created a Fusion360 plugin as shown in Figure 9 that communicates with CAMeleon and displays previews of the resulting fabrication workflows. The browser-based nature of CAMeleon makes it flexible to be integrated with modern CAD tools.
After Design. CAMeleon takes volumetric models as input, these can come from a variety of sources, one of course is a CAD model when a user is done with their basic design. CAMeleon lets them design the model agnostic to the specific fabrication workflow, just specifying its volume. They can then use CAMeleon to generate and explore different human-centered CAM workflows. They could similarly download models shared online, or use generative AI tools (such as GENIE, Meshy, TRIPO), which are specifically good at generating workflow-agnostic shape. In both cases CAMeleon allows the generic volume to be converted to practical human-centered CAM workflows.
After Fabrication. Finally, CAMeleon can serve as a bit of a re-compiler of existing models. Users may like the shape of an object they see around them, 3D scan it, and then import that geometry in CAMeleon. The result is that they can re-fabricate the same shape but chose a different fabrication process instead. Figure 10 shows an example workflow of re-fabricating a Jellcat fox.
7. extending CAMeleon
To validate the extensibility of CAMeleon, we took two distinct approaches: implement expert workflows, and capture workflows presented in research papers as human-centered CAM workflows in CAMeleon. We approached six expert craftspeople, with their own bespoke workflows, to collaborate with us in implementing their workflows as extensions of CAMeleon. After we finished the implementation, we followed-up by reflecting with them on the results. We also identified three research papers describing specific fabrication workflows and implemented those in CAMeleon. In this section we describe the resulting workflows and the process of extending CAMeleon to encapsulate these workflows.
7.1. Collaborating Craftspeople
We collaborated with six craftspeople, (Table 1 shows their demographics and background), with the aim to re-implement their workflows in CAMeleon. Beforehand we conducted informal interviews with them after showing our software to better understand their perspective and workflows.
ID | Gender | Experience | Craft | Workflow |
P1 | M | 20 years | Clay sculptor | Clay sculpture |
P2 | M | 4 years | Metal electroplating craftsman | Wire copper electroplating |
P3 | M | 20 years | Needle felter | Needle felt sculpture |
P4 | M | 30 years | Carpenter | Epoxy coating |
P5 | F | 7 years | Papier-mâché craftsman | Paper mache shaping |
P6 | F | 6 years | Cake decorator | Silicone mold making |
7.2. Informal Interviews with Craftspeople
We draw the following observations from the informal semi-structured interviews.
Choosing the Right Workflow. Participants commented on the complexity and performance of fabrication processes, and how some are more or less accessible. P6 mentioned, ”I feel laser cutting is more convenient and much faster… but if you need a very precise and smooth surface, you might need to 3D print it”. P5 emphasized the importance of convenience and simplicity: ”The first thing I think about is whether I can actually make it with the material and how easy or difficult it will be to work with. I also consider the convenience and simplicity—whether I have the patience to finish the task.”
Four of the six craftspeople serve as educators in maker spaces. Their insights highlighted the value of integrating thinking around tool exposure and material understanding into the platform regarding workflow selection. P4 emphasized the importance of helping users understand the variety of tools available: ”The tools that people use are crucial. I think you’ve represented some, like laser cutting, mold making, needle felting, and styrofoam cutting… many students have no idea what tools they can use to shape or fabricate.” P3 appreciated how CAMeleonbalances tool affordances with material constraints: ”First and foremost, it’s about understanding the affordances and constraints of the tool and material we’re using… For example, if you’re using an Ultimaker, you’re limited by the build size… If a student or someone new to 3D printing starts from there, they won’t expect to print something 8 feet long.”
Experiential Learning and Creative Exploration for Entry-Level Makers. When discussing how to effectively support entry-level makers in learning workflows and sharing knowledge, P5 shared her personal experience: ”Initially, I would see others doing it online and want to try it myself… then I would integrate my own ideas and start exploring.” P3 also emphasized the value of exploration: ”It’s about getting the students or the novice maker to recognize there’s value in that investigation… and therefore you need to really engage that process without this obsession or focus on the final product.”
P2 highlights the value of exploring options before committing: ”[entry level makers, red.] lack the information to know how to make their required object, what kind of material or structure to use. And then I also think that you don’t just need to know how to make it the way you want. You need to know all the other ways it can be done so that you can decide which way is best.” Together, these insights highlight that entry-level makers benefit from experiential learning, which forms the foundation for future creative problem-solving.
Combining CAD with craft workflow. The integration of CAD into crafts can lead to remarkable outcomes. P1 highlights the efficiency and resource-saving benefits that digital fabrication brings to the creative process,”I do think that properly integrating that digital technology into traditional craftsmanship is capable of producing some really amazing stuff. If you’re very comfortable and efficient at producing 3D models, it’s gonna save you a lot more time.” P3 suggested that basic craftsmanship skills provide a deeper appreciation of materials, and tactile experience: ”there’s this preference or obsession around digital fabrication… I would require some basic shop training before even getting to digital fabrication because then you would have at least handled a block of wood… just feeling the terror of a power saw and a table saw…” P4 reinforces the synergy between digital and physical crafting, highlighting how combining both realms opens new creative possibilities: ”Digital fabrication and craft… are two different realms… you’re modeling something… in CAD… and the other, you’re crafting something with your hands… It’s really interesting when you combine those two different things… I like that about your program because that’s what it does… it helps bridge the gap between physical creating and crafting and the digital design.”
7.3. Implementing Craft Workflows
Here we lay out the process of re-implementing the workflows of our collaborating craftspeople in CAMeleon. We specifically discuss what aspects of the foundational workflows we managed to re-use in these workflows, highlighting valuable reusable modules. Figure 11 shows the reuse of modules of the original foundational workflows in creating the six additional craft workflows.
While the re-use of existing modules made up the bulk of the code, the largest amount of time went into creating good instructions for the human-centered CAM workflows. This can typically not be reused from existing practice. Our collaborators had most of the required documentation readily available because some of them share their workflows online on platforms like Instructables. But materials existed in various levels of quality, and sometimes they tend to skip steps. We therefore re-recorded a series of interactive steps for each of the workflows. For practical extension, we added a feature in CAMeleon to link to a video or external page instead of reproducing all steps for each workflow. This reduces the barrier to entry for people with bespoke workflows.
Figure 12 presents the re-implemented craft workflows in CAMeleon, to demonstrate that these work, (b) we exported and (c) fabricated a Stanford Bunny with each of these workflows. From left to right, the processes are Wire copper electroplating, Clay Sculpting, Silicone Molding, Needle Felting, Epoxy Coating, and Paper Mache Shaping.
After implementation, we validated these workflows by showing them to the craftspeople and tuning basic parameters in the design based on their experience. The final workflows will be released as standard modules in CAMeleon with the craftspeople credited for their input.
7.4. Re-implementation of Research Workflows
To demonstrate extensibility of CAMeleon beyond classical craft workflows, we re-implemented three workflows we found in research papers. Again each of them reusing modules from our foundational workflows. With our template, we were able to quickly adapt workflows of Escape Loom (Deshpande et al., 2021), StackMold (Valkeneers et al., 2019), and Bamboo Agents (Gao et al., 2023) by adjusting parameters in key modules such as the slice generator, process convex hull, mesh intersection line, and export tools (STL, SVG). The flexibility and modularity of our system allowed for efficient adaptation with minimal changes, showcasing how the re-usability of foundational modules streamlines the integration of diverse digital fabrication workflows.
However, during the implementation process, we discovered that while the software behind these workflows excels at handling geometric calculations, it falls short in accurately recording the full craft process. Although some steps have been simplified through digital tools, the workflows are still incomplete in capturing the nuances of hands-on craftsmanship. Key aspects of fabrication, which also rely on experiential knowledge, remain outside the scope of what these tools can document. While they help reduce certain challenges, they cannot fully represent the entire craft workflow, leaving gaps in the process that only skilled craftsmen can bridge.
8. USER STUDY
To evaluate the usability of an end-to-end experience with CAMeleon, we conduct a user study.
Participants We recruited 12 participants from our university (number of female students = 8, age mean = 24.66 years, SD = 1.66). Most participants (9) reported having beginner proficiency with CAD tools.
Task. Participants were tasked to fabricate a physical model from a digital design created with Luma AI (a generative AI tool). After importing their AI-generated 3D model into CAMeleon, they explored different ways of fabricating their object and used the respective human-centered CAM workflow to fabricate their object.
Metrics. We measured to general usability experience of using CAMeleon with questions from the standard SUS survey (Brooke, 1996). This survey was presented to participants after they exported their model from CAMeleon. We furthermore conducted semi-structured interviews a day later after participants fabricated their final object.
Procedure. A member of the research team started each session by outlining the research objectives and introducing CAMeleon. This introduction took approximately 5 minutes. Participants were given 10-minute to outline what they intend to fabricate. Participants generated their target model and used CAMeleon to explore and determine their target workflow. When done, they answer a short questionnaire aimed at capturing their initial reactions and usability experiences with CAMeleon. Researchers used the exported machine-instructions to fabricate the parts over night, the following day, participants returned to manually assemble the parts using their chosen human-centered CAM workflow with interactive guidance from CAMeleon.We concluded with a 15-minute semi-structured interview to gather more detailed feedback about their interactions with CAMeleon and its effectiveness in assisting them from digital model to physical assembly.
8.1. Results
All participants succeeded in finding their desired workflow and built successfully, Figure 14 shows the resulting physical objects.
Figure 15, shows the results of the survey. Most participants found the software easy to use (10/12 gave it a score of more than 3/5), well-integrated (12/12 rated ¿3, and 8 gave it a 5/5), and not overly complex (11/12 rated this 1, 1/12 rated it 2). 10/12 participants indicated they would like to use this software frequently (2 gave it a neutral 3). Similarly, ease of use was rated positively by 10 participants and 2 gave it a neutral score.Most participants believed that people would quickly learn how to use the software (10 gave it a 5). The total mean SUS score amounted to 88.3.
On the more task-specific questions, 11/12 participants indicated the fabricated object met their expectations one was neutral on this. 8 participants indicated that they were confident in deciding their workflow, 2 gave this a negative score (2/5) and 2 were neutral. All participants indicated that they thought it was easy to change their workflow after seeing initial results and explored alternative processes.
8.2. Discussion
Positive User Experience Most participants found the software simple and easy to use, allowing those with no prior experience to get started smoothly. P1 highlighted that ”the workflow was nice, so I know what I’m doing, and how many steps are there,” appreciating the clear step-by-step guidance. P9 echoed this sentiment, stating, ”It is very direct. There is no complexity involved in it.” Similarly, P10 noted that ”the UI/UX design is pretty friendly to people like me,” even though they had little experience in fabrication. P6 also emphasized, ”I really love the detailed steps… because I don’t have previous (fabrication) experience. These steps helped me understand what to do, and made the whole process easy to follow.” These responses suggest that the software is well-designed for novice users, providing an accessible and straightforward navigation experience.
Workflow Comparison Users highlighted the visualization of workflows and comparison features as key strengths, helping them make informed decisions throughout the process. P2 noted, “I just want to see how it turns out, how it showed up, yeah, it’s more like aesthetic and a visualization,” emphasizing the importance of visual feedback in decision-making. Similarly, P5 shared, “So when I was doing that, it helped me to realize, okay, this is something that I really want to do. Maybe it was showing me a 3D model, which might be very close to how things would look like after. So it helped a lot.” P9 found the drag-and-drop comparison feature especially useful, stating, “I like how I can drag and drop the two or more workflows and, you know, compare them. I don’t think I am going to change any of it. It’s very self-explanatory and very distinguished software to use,” highlighting how this feature made it easier to select the desired production method. P8 echoed this sentiment, explaining that “the direct comparison of the workflow. It’s definitely helpful for me to find, like, finalize my choice.” Overall, participants appreciated the comparison feature for helping them understand the differences between various options, with visualization playing a critical role in decision-making.
Human-Centered CAM workflows Nearly all participants praised the step-by-step guidance, emphasizing how it helped them understand and follow the workflows. P10 highlighted, “I don’t have any previous experience, you know, having this. But the point is, you have the example for me, so which makes it super clear and explicit. It’s no matter the guidance on the software or the sample that you’re prepared for the users to, you know, anticipate what kind of results we really get.” Both P1 and P11 similarly appreciated this feature. P1 remarked, “This kind of handmade guidance. It’s usable. Yeah, it was super useful because I didn’t know what to do after.” Despite many users having limited or no prior experience in fabrication/craft, the detailed guidance provided by the software boosted their confidence in completing their projects.
Feature Requests: Expanding Support for Advanced Users Several participants noted that adding more advanced features could improve the experience for those with more expertise. One common suggestion was to include difficulty ratings and clearer pros and cons comparisons between workflows. For instance, P7 commented, “Maybe you could have a scale, like, ‘we advise you to use this one because it might be the easiest for what you want to achieve.” In the future, incorporating more advanced geometric analysis could further enhance the decision-making process for users. Similarly, P11 proposed using a star rating system or simple descriptions like “this workflow is suitable for beginners” or “this one saves time” to guide decision-making more effectively.
Several users highlighted the value of integrating generative AI functionality to simplify design and workflow selection. P6 suggested, “If you could embed that like a generative model in your software… that would be really helpful.” Incorporating AI-driven features or a chat-bot that recommends workflows based on user input could streamline the user experience, particularly for those who lack expertise in 3D modeling or fabrication.
9. CONCLUSION AND FUTURE WORK
We have presented CAMeleon an interface to explore and compare workflows for computational fabrication. We furthermore introduce the idea of human-centered CAM workflows as synergistic workflows situated between classical CAM and manual craft processes. We have evaluated the usability and utility of CAMeleon with 12 participants in our study, demonstrating a variety of workflows on volumetric models they generated. We validated the ability to extend CAMeleon with new workflows, by collaborating with six craftspeople and implementing their workflows as human-centered CAM workflows in CAMeleon. We see CAMeleon as a small step towards a future where fabrication users share and explore workflows and processes, instead of specific models for fabrication. Similar to how recipe books for cooking serve as a starting point for inspiration instead of an endpoint to replicate the exact same dish.
Moving forwards we plan to further improve the extensibility of CAMeleon, now we have identified a series of commonly reusable modules and workflows, it would be interesting to create a visual block-based programming environment where craftspeople without coding experience can remix new workflows. We furthermore plan to distribute CAMeleon in local fabrication and craft communities to study it’s longer-term usage and adoption. And following recommendations from the study, integrating AI to generate models and browse workflows within CAMeleon could streamline the overall experience.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.References
- (1)
- Goo (2010) 2010. Science, policy, and stakeholders in water management: an integrated approach to river basin management. Earthscan, London; Washington, DC.
- Alderighi et al. (2018) Thomas Alderighi, Luigi Malomo, Daniela Giorgi, Nico Pietroni, Bernd Bickel, and Paolo Cignoni. 2018. Metamolds: computational design of silicone molds. ACM Transactions on Graphics 37, 4 (Aug. 2018), 1–13. https://doi.org/10.1145/3197517.3201381
- Baudisch et al. (2017) Patrick Baudisch, Stefanie Mueller, and others. 2017. Personal fabrication. Foundations and Trends® in Human–Computer Interaction 10, 3–4 (2017), 165–293. Publisher: Now Publishers, Inc..
- Baudisch et al. (2023) Patrick Baudisch, Stefanie Mueller, Thijs Roumen, Pedro Lopes, Robert Kovacs, Jotaro Shigeyama, Shohei Katakura, Muhammad Abdullah, Conrad Lempert, Martin Taraz, and Lukas Rambold. 2023. Pushing Fabrication Research past the Makers. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–5. https://doi.org/10.1145/3544549.3583922
- Baudisch et al. (2019) Patrick Baudisch, Arthur Silber, Yannis Kommana, Milan Gruner, Ludwig Wall, Kevin Reuss, Lukas Heilman, Robert Kovacs, Daniel Rechlitz, and Thijs Roumen. 2019. Kyub: A 3D Editor for Modeling Sturdy Laser-Cut Objects. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–12. https://doi.org/10.1145/3290605.3300796
- Bourgault et al. (2023) Samuelle Bourgault, Pilar Wiley, Avi Farber, and Jennifer Jacobs. 2023. CoilCAM: Enabling Parametric Design for Clay 3D Printing Through an Action-Oriented Toolpath Programming System. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–16. https://doi.org/10.1145/3544548.3580745
- Brooke (1996) john Brooke. 1996. SUS: A ’Quick and Dirty’ Usability Scale. In Usability Evaluation In Industry. CRC Press. Num Pages: 6.
- Derhamy et al. (2018) Hasan Derhamy, Mattias Andersson, Jens Eliasson, and Jerker Delsing. 2018. Workflow management for edge driven manufacturing systems. In 2018 IEEE Industrial Cyber-Physical Systems (ICPS). IEEE, St. Petersburg, 774–779. https://doi.org/10.1109/ICPHYS.2018.8390805
- Deshpande et al. (2021) Himani Deshpande, Haruki Takahashi, and Jeeeun Kim. 2021. EscapeLoom: Fabricating New Affordances for Hand Weaving. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–13. https://doi.org/10.1145/3411764.3445600
- Dumas et al. (2014) Jérémie Dumas, Jean Hergel, and Sylvain Lefebvre. 2014. Bridging the gap: automated steady scaffoldings for 3D printing. ACM Transactions on Graphics 33, 4 (July 2014), 1–10. https://doi.org/10.1145/2601097.2601153
- Fossdal et al. (2023) Frikk H Fossdal, Vinh Nguyen, Rogardt Heldal, Corie L. Cobb, and Nadya Peek. 2023. Vespidae: A Programming Framework for Developing Digital Fabrication Workflows. In Proceedings of the 2023 ACM Designing Interactive Systems Conference. ACM, Pittsburgh PA USA, 2034–2049. https://doi.org/10.1145/3563657.3596106
- Gao et al. (2023) Peizhong Gao, Tanhao Gao, Yanbin Yang, Zhenyuan Liu, Jianyu Shi, and Jin Li. 2023. Bamboo Agents: Exploring the Potentiality of Digital Craft by Decoding and Recoding Process. In Proceedings of the Seventeenth International Conference on Tangible, Embedded, and Embodied Interaction. ACM, Warsaw Poland, 1–13. https://doi.org/10.1145/3569009.3572746
- Hartikainen et al. (2024) Heidi Hartikainen, Leena Ventä-Olkkonen, Marta Cortés Orduña, Ivan Sanchez Milara, Marjukka Käsmä, and Leena Kuure. 2024. Making The Future School: An Analysis of Teens’ Collaborative Digital Fabrication Project. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 91:1–91:37. https://doi.org/10.1145/3637368
- Higgins et al. (2022) Erin Higgins, William Berkley Easley, Karen L. Gordes, Amy Hurst, and Foad Hamidi. 2022. Creating 3D Printed Assistive Technology Through Design Shortcuts: Leveraging Digital Fabrication Services to Incorporate 3D Printing into the Physical Therapy Classroom: Leveraging Digital Fabrication Services to Incorporate 3D Printing into the Physical Therapy Classroom. In Proceedings of the 24th International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’22). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3517428.3544816
- Hurst and Tobias (2011) Amy Hurst and Jasmine Tobias. 2011. Empowering individuals with do-it-yourself assistive technology. In The proceedings of the 13th international ACM SIGACCESS conference on Computers and accessibility. 11–18.
- Iarussi et al. (2015) Emmanuel Iarussi, Wilmot Li, and Adrien Bousseau. 2015. WrapIt: computer-assisted crafting of wire wrapped jewelry. ACM Transactions on Graphics 34, 6 (Nov. 2015), 1–8. https://doi.org/10.1145/2816795.2818118
- McCrae et al. (2014a) James McCrae, Nobuyuki Umetani, and Karan Singh. 2014a. FlatFitFab: interactive modeling with planar sections. In Proceedings of the 27th annual ACM symposium on User interface software and technology. ACM, Honolulu Hawaii USA, 13–22. https://doi.org/10.1145/2642918.2647388
- McCrae et al. (2014b) James McCrae, Nobuyuki Umetani, and Karan Singh. 2014b. FlatFitFab: interactive modeling with planar sections. In Proceedings of the 27th annual ACM symposium on User interface software and technology. ACM, New York, NY, USA, 13–22.
- Mori and Igarashi (2007) Yuki Mori and Takeo Igarashi. 2007. Plushie: an interactive design system for plush toys. ACM Transactions on Graphics 26, 3 (July 2007), 45. https://doi.org/10.1145/1276377.1276433
- Nakashima et al. (2018) Kazutaka Nakashima, Thomas Auzinger, Emmanuel Iarussi, Ran Zhang, Takeo Igarashi, and Bernd Bickel. 2018. CoreCavity: interactive shell decomposition for fabrication with two-piece rigid molds. ACM Transactions on Graphics 37, 4 (Aug. 2018), 1–13. https://doi.org/10.1145/3197517.3201341
- Parry-Hill et al. (2017) Jeremiah Parry-Hill, Patrick C. Shih, Jennifer Mankoff, and Daniel Ashbrook. 2017. Understanding Volunteer AT Fabricators: Opportunities and Challenges in DIY-AT for Others in e-NABLE. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (CHI ’17). Association for Computing Machinery, New York, NY, USA, 6184–6194. https://doi.org/10.1145/3025453.3026045 event-place: Denver, Colorado, USA.
- Rai et al. (2009) Sudhendu Rai, Charles B. Duke, Vaughn Lowe, Cyndi Quan-Trotter, and Thomas Scheermesser. 2009. LDP Lean Document Production—O.R.-Enhanced Productivity Improvements for the Printing Industry. Interfaces 39, 1 (Feb. 2009), 69–90. https://doi.org/10.1287/inte.1080.0413
- Roumen et al. (2022) Thijs Roumen, Ingo Apel, Thomas Kern, Martin Taraz, Ritesh Sharma, Ole Schlueter, Jeffrey Johnson, Dominik Meier, Conrad Lempert, and Patrick Baudisch. 2022. Structure-Preserving Editing of Plates and Volumes for Laser Cutting. In Proceedings of the 7th Annual ACM Symposium on Computational Fabrication. ACM, Seattle WA USA, 1–12. https://doi.org/10.1145/3559400.3561996
- Schiff and Bates (2010) Gordon D. Schiff and David W. Bates. 2010. Can Electronic Clinical Documentation Help Prevent Diagnostic Errors? New England Journal of Medicine 362, 12 (March 2010), 1066–1069. https://doi.org/10.1056/NEJMp0911734
- Schmidt and Umetani (2014) Ryan Schmidt and Nobuyuki Umetani. 2014. Branching support structures for 3D printing. In ACM SIGGRAPH 2014 Studio. ACM, Vancouver Canada, 1–1. https://doi.org/10.1145/2619195.2656293
- Schwartzburg and Pauly (2013) Yuliy Schwartzburg and Mark Pauly. 2013. Fabrication‐aware Design with Intersecting Planar Pieces. Computer Graphics Forum 32, 2pt3 (May 2013), 317–326. https://doi.org/10.1111/cgf.12051
- Stava et al. (2012) Ondrej Stava, Juraj Vanek, Bedrich Benes, Nathan Carr, and Radomír Měch. 2012. Stress relief: improving structural strength of 3D printable objects. ACM Transactions on Graphics 31, 4 (Aug. 2012), 1–11. https://doi.org/10.1145/2185520.2185544
- Subbaraman and Peek (2022) Blair Subbaraman and Nadya Peek. 2022. p5.fab: Direct Control of Digital Fabrication Machines from a Creative Coding Environment. In Designing Interactive Systems Conference. ACM, Virtual Event Australia, 1148–1161. https://doi.org/10.1145/3532106.3533496
- Tibeme et al. (2018) Bernis Tibeme, Hossain Shahriar, and Chi Zhang. 2018. Process Mining Algorithms for Clinical Workflow Analysis. In SoutheastCon 2018. IEEE, St. Petersburg, FL, 1–6. https://doi.org/10.1109/SECON.2018.8479176
- Tojo et al. (2024) Kenji Tojo, Ariel Shamir, Bernd Bickel, and Nobuyuki Umetani. 2024. Fabricable 3D Wire Art. In Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24. ACM, Denver CO USA, 1–11. https://doi.org/10.1145/3641519.3657453
- Tran O’Leary et al. (2023) Jasper Tran O’Leary, Gabrielle Benabdallah, and Nadya Peek. 2023. Imprimer: Computational Notebooks for CNC Milling. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–15. https://doi.org/10.1145/3544548.3581334
- Tran O’Leary et al. (2024) Jasper Tran O’Leary, Thrisha Ramesh, Octi Zhang, and Nadya Peek. 2024. Tandem: Reproducible Digital Fabrication Workflows as Multimodal Programs. In Proceedings of the CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–16. https://doi.org/10.1145/3613904.3642751
- Tseng and Resnick (2014) Tiffany Tseng and Mitchel Resnick. 2014. Product versus process: representing and appropriating DIY projects online. In Proceedings of the 2014 conference on Designing interactive systems. ACM, Vancouver BC Canada, 425–428. https://doi.org/10.1145/2598510.2598540
- Tseng and Tsai (2015) Tiffany Tseng and Geoff Tsai. 2015. Process Products: Capturing Design Iteration with Digital Fabrication. In Proceedings of the Ninth International Conference on Tangible, Embedded, and Embodied Interaction. ACM, Stanford California USA, 631–636. https://doi.org/10.1145/2677199.2687891
- Valkeneers et al. (2019) Tom Valkeneers, Danny Leen, Daniel Ashbrook, and Raf Ramakers. 2019. StackMold: Rapid Prototyping of Functional Multi-Material Objects with Selective Levels of Surface Details. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology. ACM, New Orleans LA USA, 687–699. https://doi.org/10.1145/3332165.3347915
- Wall et al. (2021) Ludwig Wilhelm Wall, Alec Jacobson, Daniel Vogel, and Oliver Schneider. 2021. Scrappy: Using Scrap Material as Infill to Make Fabrication More Sustainable. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–12. https://doi.org/10.1145/3411764.3445187
- Zhao et al. (2018) Jian Zhao, Michael Glueck, Petra Isenberg, Fanny Chevalier, and Azam Khan. 2018. Supporting Handoff in Asynchronous Collaborative Sensemaking Using Knowledge-Transfer Graphs. IEEE Transactions on Visualization and Computer Graphics 24, 1 (Jan. 2018), 340–350. https://doi.org/10.1109/TVCG.2017.2745279
Appendix A Workflows Sources
Here is the list of workflows sources with links to their instruction pages.