### v2

# CAMeleon: Interactively Exploring Fabrication Workflows in CAD

Abstract.
Designers of physical objects make assumptions on the material and fabrication workflow early in the design process. Recovering from bad assumptions is hard, because the design and resulting CAD model are locked-in to those assumptions. We present CAMeleon, a software tool to interactively explore and compare fabrication workflows at the end of the design process.
CAMeleon’s modular architecture allows users to execute their design with different workflows, and preview results. Users can explore alternative workflows. CAMeleon’s architecture is extensible with new workflows, increasing the scope of workflows available.
Based on a survey of 150 fabrication workflows shared online, we implemented five fabrication representative workflows in CAMeleon. To validate CAMeleon, we collaborated with six craftspeople, replicating their workflows to gain insights into their practices. Additionally, we demonstrated CAMeleon’s extensibility by implementing workflows from three research papers and reflecting on the extension process. Our design workshop (16 participants) highlights how working with CAMeleon unlocks a creative attitude of exploring workflows and opened up participants to workflows they had not previously imagined. We conclude that tools like CAMeleon have the potential to drastically expand designer’s ability to experiment with new materials and fabrication workflows.
1. INTRODUCTION
Advances in computer-aided design (CAD) software and computational fabrication (e.g., additive manufacturing, subtractive fabrication, and wire forming) have enabled an increasingly broad set of people to engage in the design of physical objects. This trend towards personal fabrication (Baudisch et al., 2017) enabled marginalized communities to create custom assistive technology (Hurst and Tobias, 2011; Parry-Hill et al., 2017; Higgins et al., 2022), allowed high-schools to make fabrication part of their curriculum (Baudisch et al., 2019; Roumen et al., 2022; Hartikainen et al., 2024), and empowered several exciting hardware start-ups (e.g., Prusa3D, BambuLab, and XTOOL). As this trend continues, some envision fabrication machines to become as ubiquitous as kitchen appliances (Baudisch et al., 2023).
However, designing physical objects presents challenges, particularly for people who are not well-versed in fabrication workflows. Before even starting to design objects, designers make assumptions about the final fabrication workflow (e.g., Laser cutting, 3D Printing, Vacuum forming …)—choices that determine how they will design the structure, select the material, operate the fabrication machine, and perform the appropriate post-processing. This is challenging because they may not know what workflows are available to them in the first place, or how choosing one workflow over another has implications for the final design.
To address these challenges, HCI researchers have developed tools to assist users really good at designing with a specific process/material in mind. Tools like Kyub (Baudisch et al., 2019) and FlatFitFab (McCrae et al., 2014a) are specialized for laser cutting, Bridging the gap (Dumas et al., 2014) and Stress relief (Stava et al., 2012) for 3D printing, and Plushie (Mori and Igarashi, 2007) for textile design. These tools have made computational fabrication accessible by reducing the required upfront domain knowledge of fabrication processes and materials. The price of using such software is that designers lock themselves into one specific fabrication process from the get-go (e.g., designing for laser cutting or for 3D printing). Ironically, as designers make progress on their design, they learn more about limitations/constraints imposed by the fabrication processes they chose, while at the same time, it gets ever harder to explore alternatives. For example, for the bunny of Figure 1, it may be hard to achieve some details in the ears using wire bending which designers only find out when they get to that detail, or it may be better to laser cut an internal structure if the designer decides to fabricate at larger scale later on in the process. Recovering from these early assumptions is challenging because many of the parameters of the CAD model (wires, plates and joints, or volumes) are now specific to the initially chosen workflow.
On the other end of the spectrum, there is the industrial workflow where engineers first design in CAD and then use so-called ”Computer Aided Manufacturing” (CAM) to convert their designs to machine-specific instructions. While CAM generates instructions for different machines, it still requires a design that is optimized for that specific fabrication process. These CAM instructions are heavily machine-oriented (e.g., they convert STLs to g-code for printing) and fail to acknowledge that most fabrication workflows involve manual processes (such as assembly, polishing, curing, etc.), which in practice highly influence the decision of what workflows to use.
In this paper, we propose CAMeleon, a software tool to interactively explore craft workflows. CAMeleon, allows users to defer their final fabrication decision and explore and compare fabrication workflows until they start fabricating. It achieves this by encapsulating fabrication workflows as modular components of the interface, each taking volumetric shape as input, converting that to workflow-specific geometry, and exporting instructions for both machines and humans. We call these instructions human-centered CAM workflows. As the target volumetric shape remains unchanged, users can freely explore alternative workflows before making final decisions. The modular architecture of CAMeleon makes it easily extensible, further increasing the richness of workflows available to end-users.
We surveyed 150 fabrication workflows shared online, we classified these into categories, and from each of these categories we implemented one representative core workflow in CAMeleon. To validate CAMeleon, we collaborated with six craft experts, replicating their workflows to gain insights into their thinking. Additionally, to study CAMeleon’s extensibility, we implemented workflows from three research papers. In a design workshop with 16 participants, we found that participants use CAMeleon to develop an attitude to explore fabrication workflows, resulting in them making good decisions about their final workflow even if they had never used some of them before.
We conclude that deferring the decision of fabrication workflows, and making this decision interactive, allows designers to develop an exploratory attitude on workflows for computational fabrication. Especially designers with expertise that deviates from specific fabrication workflows, benefit from the ability to revise their decisions later on. As shape-based CAD models become more accessible, and tools like CAMeleon provide methods to capture human-centered CAM workflows, similar to recipes for cooking appliances, we hope to take a small step towards reducing the barrier to entry for many more users to start taking advantage of their ability to create a physical reality that suits their needs.
2. WALK-THROUGH: exploring workflows to fabricate a stool
Figure 2 illustrates the process of Peter creating a small stool for his child using CAMeleon.
Peter is a designer who has some experience with 3D printing but not much with other fabrication workflows. Initially, he attempted to design a stool model for 3D printing following online tutorials—adjusting parameters such as thickening walls, reducing overhangs, and modifying infill settings to meet 3D printing requirements. However, after completing the print, Peter realized that his available 3D printing material lacks the inherent strength needed for the stool, making it unsuitable. Additionally, the surface finish did not meet his expectations.
Unsure how to proceed, Peter searches for alternative workflows. Platforms like Instructables provide useful tutorials and step-by-step instructions for specific projects, but Peter finds it difficult to adapt his design to those workflows. Furthermore, while many workflows seem promising, he struggles to visualize how the final fabricated results would look for his model. To compare workflows, he has to navigate between multiple web pages, gathering fragmented information on fabrication methods and repeatedly adjusting his design. This process is time-consuming and impractical.
Frustrated with this fragmented process, Peter turns to CAMeleon, which runs straight in his browser. CAMeleon allows Peter to explore and compare various workflows rendered as visualization of the expected outcomes.
As shown in Figure 2a, Peter begins by downloading a 3D volumetric model (here an STL file) from Thingiverse, a repository of 3D models. He uploads this model into CAMeleon. (b) Unsure of which workflow best meets his needs, Peter starts by exploring the available options. He browses through some options and becomes familiar with the various workflows and the machines required for fabrication, displayed as icons beneath each workflow. Peter has access to a laser cutter, a 3D printer, and a wire bender, which allows him to consider different approaches for creating the stool. Given the need for a strong material with a smooth finish for his child’s safety, Peter explores and compares workflows that meet his requirements, he drags the ’Mold making,’ ’Wire mesh sculpture,’ ’Thermal Wire Cutting for Foam Fabrication’, and ’Epoxy coating’ workflows to the view-port. After comparing the resulting visualizations, as well as the manual and machine-based manufacturing processes involved, Peter decides that the ’Epoxy coating’ workflow is the most suitable. (c) He changed the layer height to 3 mm to fit the material he has in store. Once he is satisfied with his design, Peter exports the machine-instruction files (here SVG for laser cutting).
After exporting, CAMeleon assists Peter step by step in executing the manual steps to build this stool. Although the guidance serves as a general description, CAMeleon provides detailed instructions and parameters to ensure that users understand each step more deeply and can replicate the process effectively.
Figure 3 illustrates the interactive human-centered CAM workflow. (a) Peter adjusts the provided standard guidance to fit his specific parameters and performs the laser cutting. (b) Peter then assembles and glues the pieces to form the structure. (c) Peter completes the assembly by sanding the laser-cut structure to achieve a smooth finish. While this step might be challenging for novice makers, CAMeleon provides tutorial links to guide users through the process. (d) Finally, Peter applies a layer of epoxy resin, carefully coating the entire structure for a sealed and polished finish.
3. CONTRIBUTIONS, BENEFITS, AND LIMITATIONS
This paper makes three core contributions:
(1) The development of an interactive design tool, CAMeleon, that allows users to explore and compare fabrication workflows. We discuss the implementation and design decisions that went into the development of CAMeleon and open source the code for others to extend.
(2) The concept of human-centered CAM workflows to include manual operations to capture realistic fabrication processes/workflows beyond mere machine instructions.
(3) The sample implementation of 14 fabrication workflows, including those derived from collaborations with craftsmen and recent research papers, to provide interactive fabrication guidance.
We evaluate these contributions with expert evaluation to include new workflows, and a design workshop to study usability and the role of CAMeleon in a collaborative design process.
Limitations include that not every model is suitable for all workflows; occasionally, there are size or other technical requirements specific to a process that prevent conversion. Processes in CAMeleon include warnings to users about such requirements (e.g., when laser cutting: ”features in this model are smaller than the chosen material thickness, you may experience a loss in fidelity”), and its interactive preview provides users with a sense of the expected outcome to sanity-check results before converging on a given workflow. In its current form, CAMeleon assumes one fabrication workflow for the entire model, while expert users can break down models into different sub-assemblies to then use different workflows (as evidenced by a participant in our workshop), there is currently no inherent support for this. Another limitation is that, in its current state, implementing new workflows requires knowledge of CAMeleon’s code base and Three.JS. In the future, we plan to develop an interface to design extensions for CAMeleon to make it more accessible to add new fabrication processes.
4. RELATED WORK
This work builds on research in software for personal fabrication, design for volumetric fabrication, and documenting workflows.
4.1. Software for Personal Fabrication
Modern fabrication processes, such as 3D printing, laser cutting, and CNC milling, have empowered users to personally fabricate custom objects (Baudisch et al., 2017). Software tools to support such fabrication processes have lowered the barrier to entry for novice users to create high-quality products. They typically support users with domain knowledge in a specific fabrication process. CAD tools for fabrication such as kyub (Baudisch et al., 2019), FlatFab (McCrae et al., 2014b) are specifically designed for one fabrication process, here laser cutting. Other tools like CoilCAM (Bourgault et al., 2023) focus on clay printing, or Imprimer (Tran O’Leary et al., 2023) for CNC milling take a more creative approach to allow users to experiment with these processes. They typically use a form of computational notebooks, such as p5.fab(Subbaraman and Peek, 2022) which allows exploration of processes across different techniques.
4.2. Design for Volumetric Fabrication
Fabricating volumetric shapes often relies on various support structures, such as molds and jigs. Each serves different purposes, such as facilitating shaping (e.g., plaster molds in ceramics, templates in woodworking) and reinforcing structures (e.g., rebar molds in construction, wooden frameworks in shipbuilding) or ensuring dimensional accuracy.
Tools like WrapIt facilitate manual wire bending for jewelry by generating jigs from 2D patterns, enabling users to easily hand-bend wires into desired shapes (Iarussi et al., 2015). Fabricable 3D Wire Art generates jigs based on 3D models and images, enabling more complex designs (Tojo et al., 2024). In laser cutting, StackMold aids designers by generating molds for prototyping multi-material and multi-colored objects embedded with electronics, streamlining the fabrication process (Valkeneers et al., 2019). Schwartzburg et al.’s Intersecting Planar Pieces provides techniques for creating structures through precise cuts and component assembly, enhancing the versatility of 2D subtractive fabrication (Schwartzburg and Pauly, 2013). In the field of 3D printing, Schmidt et al. developed automated support generation systems that stabilize complex prints during 3d printing (Schmidt and Umetani, 2014). Innovations like CoreCavity offer interactive shell decomposition methods to generate two-piece rigid molds for diverse materials (Nakashima et al., 2018), while Metamolds produce 3D-printed containers designed for creating silicone molds, expanding the material possibilities for designers (Alderighi et al., 2018). Diverging from traditional support methods, Scrappy introduces a novel approach by utilizing scrap materials as infill for 3D prints, promoting sustainability and resource efficiency in fabrication processes (Wall et al., 2021).
These tools effectively assist in shaping during craft processes or converting existing models into complex structures, thereby lowering the technical barriers and making these processes more accessible. Building on this background, our goal is to create a tool whih allows users to explore and compare such workflows, allowing more people to access different techniques. This platform aims to bridge the gap between novice users and craftspeople, serving as both a resource for learning about craft techniques and an educational tool for sharing craft knowledge.
4.3. Documenting Workflows
In disciplines such as industrial manufacturing (Derhamy et al., 2018), and business process management (Goo, 2010), there has been a long standing focus on documenting workflows and processes. Such documentation has been instrumental to increase efficiency (Rai et al., 2009), detect and minimize errors (Schiff and Bates, 2010), and support continuous process improvement (Tibeme et al., 2018). It furthermore facilitates the transfer of knowledge and ideas across teams or enterprises (Zhao et al., 2018). In craft communities such workflow knowledge often is not transferred from one to another and fails to benefit from similar distribution due to unpredictable material behaviors (Ingold, 2010) and domain-specific tacit knowledge (Sennett, 2008).
To tackle this challenge, researchers have started integrating multiple fabrication workflows into unified platforms, significantly advancing the field of digital fabrication. For example, Vespidae offers flexibility in controlling fabrication machines, allowing them to manipulate key parameters. It features modular components that users can freely combine to form new code definitions, facilitating a deeper exploration of the design space (Fossdal et al., 2023). Similarly, Tandem addresses the complexity of experimental digital fabrication workflows, successfully capturing and recording workflows across multiple software tools, including machine tuning and manual interventions (Tran O’Leary et al., 2024). Research such as Process Products captures iterative processes by embedding this information directly into objects, which helps designers communicate design evolution and aids manufacturers in understanding tool usage (Tseng and Tsai, 2015). Similarly, Product versus Process explores how design documentation is created and used by authors and readers on platforms like Instructables111https://www.instructables.com/ (Tseng and Resnick, 2014).
On the hobbyist side of the spectrum, online communities such as Instructables, Dorkbot222https://dorkbot.org/ and ravelry333https://www.ravelry.com/, serve as knowledge-sharing hubs where makers showcase their projects. These platforms allow users to share detailed documentation of fabrication processes, including the use of machines, software, materials, and techniques, while engaging in discussions with other makers to exchange tips and experiences. On the other hand, most programming tools for fabrication primarily focus on the digital fabrication part of the process, without delving into how users can approach subsequent craft techniques or enhance their projects through hands-on methods. P5.fab (Subbaraman and Peek, 2022) make this case by opening up CAM workflows to more creative manipulation. For example, tools like WrapIt offer strong support in generating jigs for wire bending (Iarussi et al., 2015), EscapeLoom enhances the design space for weaving by building support structures (Deshpande et al., 2021), and Metamolds automatically defines the cuts needed to extract cast objects from silicone molds (Alderighi et al., 2018). While these tools excel in their respective areas, they could be complemented with resources that provide users guidance on the manual techniques required in these crafts.
5. Surveying craft workflows
The goal of CAMeleon is to let users explore fabrication workflows. To better understand the space of existing workflows, we explored workflows shared in online communities. We reviewed 150 winning projects from the ’Workshop’ category on Instructables444https://www.instructables.com (sorted by ’I Made Its’). We excluded workflows centered on microprocessor-based products, purely 3D printed items without any additional workflow, and those focused solely on polishing, to maintain an emphasis on fabrication workflows. This left us with 55 projects. The appendix contains a list of all workflows used for this process with links to their instruction pages. We proceeded to cluster these into categories to better understand the design space. In Figure 4, the different processes are shown and divided into the following categories: Mold Casting, Stacked Slice Construction, Interlocking Structure, Guide Structure, and Wire Forming. Two workflows did not fit this categorization; one of them was creating shapes with epoxy resin, and the other was building boats from wooden spars and frames.
To start off with a representative subset of fabrication processes based on this classification, we implemented one representative sample of each category in CAMeleon, we call these representative workflows. Figure 5 shows an overview of these at the example of the Stanford Bunny:(a) Thermal Wire Cutting Foam, (b) Mold Casting, (c) Laser-Cut Interlocking Structures, (d) Wire Mesh Fabrication, and (e) Stacked Slice Construction. The variety of these processes not only gives users five interesting workflows to start out with, but they also serve as good templates as building blocks for others to produce workflows.
6. CAMeleon
CAMeleon is a platform to explore and compare fabrication workflows. CAMeleon aims to support the process of computational fabrication and human-centered CAM workflows. We have the following design goals:
Shared and Reusable. CAMeleon’s code base should be modular by exposing workflow functions as reusable modules to enable extensibility.
Accessible Interface Design. The goal is to give users all the information needed to compare and explore workflows in a nutshell. Therefore, we focus on ease of use and good overviews. The hope is that designers at various levels of expertise are able to come to a useful comparison/decision.
Inclusive Open Platform. To bring together designers with different backgrounds, CAMeleon should not depend on existing software, and therefore should be implemented as browser tool. This furthermore ensures broad compatibility across devices and operating systems.
Compatible with Machines. To maximize support across fabrication machines, CAMeleon should support exporting to popular extensions for 3D (STL), 2D (SVG) and 1D/other (CSV) machines.
In this section, we discuss design considerations and technical details that form the basis of CAMeleon. We discuss the interface design and the implementation.
6.1. Interface Design
CAMeleon is a browser tool. Users compare workflows using the comparison tool and the ability to browse different workflows. Once a workflow is selected, the other core aspect of the interface is the interactive guides to create human-centered CAM workflows. We discuss each individually in this subsection.
6.1.1. Comparison Windows
The comparison tool (shown in Figure 6) lets users drag up to four workflows into the view-port. All workflows load the same 3D model, enabling users to compare potential outcomes side by side. In addition to visual comparison, the comparison windows also include information about the processes involved in each workflow.
Users can preview key manual steps and required machines, as outlined in the human-centered CAM workflows. This feature allows workflows to be compared not only by their visual results but also by the steps and resources they require. Users can close workflows to compare with others or free up space to focus on 1-2 workflows in more detail.
6.1.2. Workflow Browser & Filtering
CAMeleon has a workflow browser panel on the left of the interface (demonstrated in Figure 7). The most straight-forward is to scroll through the workflows, however this does not scale as the number of workflows increases. Users therefore can filter by keywords related to the specific workflow or meta-properties such as workflow criteria, machines, or materials.
Using the workflows from our survey, we identified three categories of workflow-related dimensions which are presented to users as shown in Figure 7(b).
Product-related Dimensions: Product properties such as high-temperature tolerance and the fidelity required to reproduce intricate details.
Structure-related Dimensions: The use and construction of support structures, jigs, or molds. This categorization helps users understand and manage the specific requirements and characteristics of the support structures used in their fabrication process.
Machine-related Dimensions: The choice of fabrication machinery is closely linked to the selection of fabrication workflows. These dimensions guide users in identifying and selecting workflows based on the machines required, such as 3D printers, laser cutters, or wire benders.
6.1.3. Interactive Guide for human-centered CAM workflows
As mentioned before, traditional CAM tools focus on machine-instructions, and fail to deliver instructions to humans who use the resulting artifacts and/or control the machines. We therefore propose human-centered CAM workflows which contains machine-instructions complemented with interactive guidance for users.
Figure 8 shows such guidance, users click through the steps one at a time. (a) This guide includes textual information, detailing how to assemble each component, with tips highlighted to ensure clarity. (b) Additionally, the guide includes links to external resources, such as YouTube tutorials, for in-depth guidance on the specific technique or examples of similar processes. These guides are currently not specific to the design of the user, they serve as a general explanation of the workflow, future work could explore how to leverage dynamic rendering techniques to contextualize these instructions for the model at hand.
6.1.4. Loading tool
The loading tool enables the import of STL models. However, when users adopt a model for a specific fabrication workflow outside CAMeleon during the design phase, the model may be constrained to a specific workflow (say, a structure of laser-cut plates). CAMeleon’s import recovery aims to reconstruct a volumetric representation, enabling users to explore alternative workflows again. Two reconstruction methods are supported: Voxelization with Poisson surface reconstruction and Alpha shape reconstruction, each addressing different structural characteristics.
Figure 9 illustrates the application of these two methods using the Stanford Bunny model. For a laser-cut interlocking planar structure (Figure 9a), Alpha shape reconstruction accurately recovered the inner geometry. For this wire-bending structure (Figure9 b), Voxelization with Poisson surface reconstruction produced a smooth and complete recovery of the outer surface. However, if a model is already volumetric, there is no need for pre-processing other than basic mesh repair (filling holes, completing triangles). The preview of the different imports gives users a sense of what their model will look like in CAMeleon as a result of the different pre-processing algorithms.
7. Implementation
CAMeleon is built on Node.js, Three.js (using WebGL) and Python. Three.js serves as the foundation, providing 3D rendering and geometry handling in the browser. Building on this, CAMeleon extends Three.js with CAD-specific geometry processing, contour simplification, projection generation, and visualization tools (e.g., draft angle analysis, bounding box helpers, and color-coded feedback). It supports output formats like STL, SVG, and CSV, tailored to different machines. The Python-based back-end complements these features by handling computationally intensive geometry operations, ensuring efficient processing.
CAMeleon’s workflows are implemented using a modular architecture. Each workflow leverages reusable modules to address tasks, such as mold making, foam cutting, or interlocking structure generation. By remixing these modules, developers compose new workflows as demonstrated in the ”Extending CAMeleon” section.
As shown in the Figure 10, When users load a new model, the WorkflowController calls the Model Loading for the specialized task of 3D model loading and pre-processing. Once the model is loaded, the front-end communicates with the Flask-based back-end through REST APIs. These APIs facilitate file upload and initiate geometry operations, leveraging libraries like Open3D and PyMeshLab for computationally intensive tasks such as model simplification, Alpha Shape generation, and voxelization.
7.0.1. Workflow Modules
CAMeleon modules are organized into four categories and listed in Table 1, covering the complete workflow for 3D visualization and fabrication:
Model Loading. Handles the import and initial processing of 3D models, including STL file loading and basic geometry adjustments.
Geometry Processing. Performs core computational tasks on 3D geometry, such as intersections, contour simplifications, and projections for specific workflows.
Visualization. Enhances understanding of 3D models through tools like draft angle analysis, dimension labeling, model information helpers, and color-coded feedback for manufacturing validation.
Export. Converts processed models or data into formats such as STL, SVG, or CSV for fabrication workflows.
7.0.2. Distribution
CAMeleon is available as a simple browser tool on RELEASED ON ACCEPTANCE, and the source code can be accessed via RELEASED ON ACCEPTANCE.
Category | Module | Description |
Model Loading | Model Loading | Load and process STL files, with options for centering, scaling, mesh rotation (x/y/z), and bounding box handling. |
Geometry Processing | Hollow Triangle Gen | Generate hollow triangles from vertices, faces, and normals. |
Threaded Rods Gen | Create 2D and 3D shapes for threaded rods. | |
Tube Gen | Generate tubes with specified color, opacity, and radius. | |
Interlock Joints Gen | Create interlocking joints based on points and dimensions. | |
Split Volume | Create positive and negative molds with air vents. | |
Intersect | Generate intersections and contours using meshes and planes. | |
Extrude | Generate 3D extruded shapes from lines. | |
Extrude to Direction | Extrude shapes in a specific direction with defined depth. | |
Loft | Create lofted surfaces from inner and outer contours. | |
Calc Draft Angles | Analyze draft angles with color mapping. | |
Project Mesh | Project meshes into 2D lines. | |
Simplify Geom | Simplify geometry based on distance and angle thresholds. | |
Calc Centroid | Compute the centroid of triangular vertices as a Vector3. | |
Calc Line Inters | Calculate intersection points for lines in specific planes. | |
Vertex Scaling | Scale vertices around a centroid using a scale factor. | |
Visualization | Model Info Vis | Display mesh details and properties. |
Visualize Mold | Visualize positive and negative mold parts. | |
Tutorial & Info | Provide text, images, and links for user guidance. | |
Color Legend | Show visual legend for angle ranges and other properties. | |
3D Visualization | Render meshes in 3D. | |
Geom Visualization | Visualize points and lines with legends. | |
Export | Export SVG | Save lines as SVG files. |
Export CSV | Save lines as CSV files. | |
Export STL | Save meshes as STL files. |
8. Expert Evaluation of CAMeleon with craftspeople
To validate CAMeleon’s ecological validity, we interviewed and collaborated with six expert craftspeople to replicate their bespoke workflows inCAMeleon, we reflect with them on the resulting workflows. This section outlines the resulting workflows and the process of extending CAMeleon to support these workflows.
8.1. participants
Table 2 shows the demographics and background of our collaborating expert craftspeople and the workflows they choose to integrate into CAMeleon.
ID | Gender | Experience | Craft | Workflow |
P1 | M | 20 years | Clay sculptor | Clay sculpture |
P2 | M | 4 years | Electroplating | Wire copper electroplating |
P3 | M | 20 years | Needle felter | Needle felt sculpture |
P4 | M | 30 years | Carpenter | Epoxy coating |
P5 | F | 7 years | Papier-mâché craftsman | Papier-mâché shaping |
P6 | F | 6 years | Cake decorator | Silicone mold making |
8.2. procedure
We approached this in three phases: first we showed participants CAMeleon and conducted semi-structured interviews to understand their perspective and workflows. We then proceeded by implementing their workflows in CAMeleon and follow-up with an interview to reflect and collect their feedback.
8.3. Initial Interviews
We draw the following observations from the initial interviews.
Choosing the Right Workflow. Participants commented on the complexity and performance of fabrication processes, and how some are more or less accessible. P6 mentioned, ”I feel laser cutting is more convenient and much faster… but if you need a very precise and smooth surface, you might need to 3D print it”. P5 emphasized the importance of convenience and simplicity: ”The first thing I think about is whether I can actually make it with the material and how easy or difficult it will be to work with. I also consider the convenience and simplicity—whether I have the patience to finish the task.”
Four experts also serve as educators. Their insights highlighted the value of integrating thinking around tool exposure and material understanding into the platform regarding workflow selection. P4 emphasized the importance of helping users understand the variety of tools available: ”The tools that people use are crucial. I think you’ve represented some, like laser cutting, mold making, needle felting, and styrofoam cutting… many students have no idea what tools they can use to shape or fabricate.” P3 appreciated how CAMeleon balances tool affordances with material constraints: ”First and foremost, it’s about understanding the affordances and constraints of the tool and material we’re using… For example, if you’re using an Ultimaker, you’re limited by the build size… If a student or someone new to 3D printing starts from there, they won’t expect to print something 8 feet long.”
Experiential Learning and Creative Exploration for Entry-Level Makers. When discussing how to effectively support entry-level makers in learning workflows and sharing knowledge, P5 shared her personal experience: ”Initially, I would see others doing it online and want to try it myself… then I would integrate my own ideas and start exploring.” P3 also emphasized the value of exploration: ”It’s about getting the students or the novice maker to recognize there’s value in that investigation… and therefore you need to really engage that process without this obsession or focus on the final product.”
P2 highlights the value of exploring options before committing: ”[entry level makers, red.] lack the information to know how to make their required object, what kind of material or structure to use. And then I also think that you don’t just need to know how to make it the way you want. You need to know all the other ways it can be done so that you can decide which way is best.” Together, these insights highlight that entry-level makers benefit from experiential learning, which forms the foundation for future creative problem-solving.
Combining CAD with craft workflow. The integration of CAD into crafts can lead to remarkable outcomes. P1 highlights the efficiency and resource-saving benefits that digital fabrication brings to the creative process,”I do think that properly integrating that digital technology into traditional craftsmanship is capable of producing some really amazing stuff. If you’re very comfortable and efficient at producing 3D models, it’s gonna save you a lot more time.” P3 suggested that basic craftsmanship skills provide a deeper appreciation of materials, and tactile experience: ”there’s this preference or obsession around digital fabrication… I would require some basic shop training before even getting to digital fabrication because then you would have at least handled a block of wood… just feeling the terror of a power saw and a table saw…” P4 reinforces the synergy between digital and physical crafting, highlighting how combining both realms opens new creative possibilities: ”Digital fabrication and craft… are two different realms… you’re modeling something… in CAD… and the other, you’re crafting something with your hands… It’s really interesting when you combine those two different things… I like that about your program because that’s what it does… it helps bridge the gap between physical creating and crafting and the digital design.”
8.4. Insights to support workflows in CAMeleon
The interviews provided valuable insights into how craftspeople consider workflows, support entry-level makers, and balance traditional crafting with digital fabrication.
A frequent theme was balancing dimensions like tool affordances, surface quality, and material constraints when selecting and comparing a workflow. Participants emphasized that providing clear and accessible guidance not only empowers users to navigate these considerations effectively but also instills in them a sense of confidence and capability to make informed choices about their projects.
Another key takeaway was the immense value of experiential learning for novices. Rather than simply comparing workflows, participants highlighted the need to provide users with essential information to deepen their understanding of tools and materials. This approach, which emphasizes hands-on experience, lets users quickly try different methods, accelerating their learning and supporting hands-on exploration. These discussions underscore the importance of designing learning platforms that not only convey technical skills but also cultivate an experimental mindset, encouraging users to appreciate both the ease of digital fabrication and the tactile, sometimes unpredictable, nature of traditional craft methods.
The interviews also revealed the dynamic interplay between traditional craftsmanship and digital fabrication. While digital tools such as CAD provide efficiency and precision, the tactile understanding developed through hands-on craftsmanship remains crucial. Participants emphasized how integrating these two realms opens new creative possibilities, allowing users to transition between designing in the digital space and physically crafting their projects.
8.5. Implementing Craft Workflows
We lay out the process of implementing the workflows of our collaborating craftspeople in CAMeleon. This required close collaboration with the craftspeople to understand their bespoke practices and translate them into CAMeleon worfklows.
The most time-consuming aspect was creating good instructions for human-centered CAM workflows. This can typically not be reused from existing practice. Our collaborators had most of the required documentation readily available. But documentation existed in various levels of quality, and sometimes they skip steps. We therefore recorded a series of steps for each of the workflows. For practical extension, we added a feature in CAMeleon to link to a video or external page instead of reproducing all steps for each workflow. This reduces the barrier to entry for people with bespoke workflows. A complete example of a human-centered CAM workflow for Needle Felting is provided in the appendix to illustrate the detailed steps and instructions.
Figure 11 presents the re-implemented craft workflows in CAMeleon, to demonstrate that these work, we exported and fabricated a Stanford Bunny with each of these workflows. From left to right, the processes are Wire copper electroplating, Clay Sculpting, Needle Felting, Papier-mâché Shaping Silicone Molding, and Epoxy Coating.
After implementation, we validated these workflows by showing them to the craftspeople. The workflows will be standard modules in CAMeleon with the craftspeople credited for their input.
8.6. Feedback
After seeing their workflows in CAMeleon, craftspeople highlighted the importance of including detailed guidance in Human-Centered CAM workflows, which they saw as crucial for supporting users. P5 emphasized the importance of explaining material choices and techniques, sharing insights on working with pulp and adhesives: “If you add too much glue and grind the paper too finely, it becomes unmanageable… Once it dries, you can mix the pulp and glue like clay to shape it.” Similarly, P6 offered insights on silicone mold workflows, discussing disassembly methods and material selection, such as, “Two-part molds reduce breakage risks… whereas single molds might crack more easily at sharp corners.” P4 suggested adding more detailed guidance on tool selection, such as ”If you needed to cut something, what type of saw would you use, or would you use a chisel? Or would you use a drill or jigsaw?…It has real teaching potential, showing users what’s possible and providing an estimate of how the process will flow.” Human-Centered CAM was praised for its ability to bridge computational design, digital fabrication, and physical workspaces. P1 noted that CAMeleon encourages users to think about physical structures and manufacturing during the design phase: “It allows users to integrate modeling with considerations for physical structures and fabrication.”
CAMeleon’s educational potential stood out to participants. P4: “It’s great for first-time users or people without sculptural experience.” P4 appreciated the diversity of techniques showcased—such as foam cutting or papier-mâché—stating, “You’re presenting options users might never have considered.” P3 emphasized how CAMeleon fills gaps in curricula:“The curriculum is often missing basic design and making knowledge. Your software helps guide novice makers to recognize the value of exploration.” Furthermore, P3 added: “We have advanced tools, but we’re hesitant about novice makers using them. Your software fills that gap, helping students gain both access and understanding.”
Analysis feature like the draft angle analysis received acclaim. P6: “It identifies weak areas—if there’s too much blue, the mold design might need adjustments.” P6 also suggested enhancing computational capabilities, such as adding gates to support delicate mold areas: “For example, supporting the ears of a chocolate rabbit mold can reduce breakage risks.”
9. Technical evaluation on the Extensibility of CAMeleon
To do a deeper dive into the extensibility of CAMeleon, we implemented workflows described in three research papers. Traditionally, research workflows focus on their core academic contributions, and thus can be hard to replicate. We demonstrate how the modular architecture of CAMeleon supports this process and thus creates a platform for experimentation across workflows beyond bespoke crafts.
9.1. Selected workflows
EscapeLoom (Deshpande et al., 2021), StackMold (Valkeneers et al., 2019), and Bamboo Agents (Gao et al., 2023) were selected as representative workflows due to their alignment with the overarching theme of integrating traditional crafts with digital fabrication tools. Each of these works, sourced from the ACM Digital Library under the Human-centered computing Index Terms, emphasizes interactive systems and tools that facilitate novel approaches to craft-making. The decision to replicate these workflows stems from their focus on enhancing user creativity and expanding the possibilities of fabrication through the interplay of manual craftsmanship and computational methods.
9.2. Results
Figure 12 shows how each of the workflows implemented in CAMeleon allowed exported the resulting Stanford Bunny. Each of the workflows could be built up from modules existing in CAMeleon. Figure 13 outlines this reuse on a meta-level.
In re-implementing workflows, we used the Model Loading module to import STL files and prepare models for subsequent geometry processing. Each workflow involved distinct processing steps tailored to its goals:
For EscapeLoom, organic surfaces were converted into frame structures by adding warps parallel to surface curves and columns for closed surfaces. This was achieved using the Intersect module for dividing surfaces into frames and the Tube Gen module for creating volumetric structures. Fabrication was supported through the Export STL module.
In StackMold, 3D models were divided into planar contours to create layered molds. The Intersect module sliced the geometry and dealt with the bounding box into contours, while the Simplify Points/Lines refined inner paths and outer boundaries. The Loft module connected the contours into molds, and the Extrude to Direction module transformed them into volumes for Visualization, with cutting paths generated via the Export SVG module.
For Bamboo Agents, hollow frame structures were constructed by dividing models into triangular elements using the Hollow Triangle Gen module. Dimensions were adjusted with the Vertex Scaling module, and the Loft and Extrude to Direction modules were used to get the main structure and fixed structure. Fabrication was facilitated through the Export STL module.
Across all workflows, modules such as Model Information Visualization, 3D Visualization, and Tutorial & Info provided visual feedback, interactive guidance, and iterative support, ensuring a smooth and efficient design process. Figure 13 highlights the modules reused in adapting these workflows.
During the implementation process, we discovered that while the software behind these workflows handles geometric calculations, it falls short in accurately recording the full complexity of the workflow. Workflows still require tacit knowledge of craftsmanship. Key aspects of fabrication, which also rely on experiential knowledge, remain outside the scope of what these tools can document. While they help reduce certain challenges, they cannot fully represent the entire craft workflow, leaving gaps in the process that only skilled craftsmen can bridge. Such tacit knowledge is often overlooked or not documented in research papers. Once more, this calls for better documentation of Human-Centered CAM workflows. Capturing its full complexity would require adding tutorials and learning interfaces. Such implementation, while valuable, is beyond the scope of our current implementation and remain an area for future exploration.
10. User Evaluation: a Design Workshop
To understand the practical utility of CAMeleon for designers, and to observe its role in a design process, we conducted a design workshop in which teams of designers design and fabricate objects of their choice. Our objective was to study what role CAMeleon plays and how/whether it facilitates the intended interactive exploration of fabrication workflows.
Participants. We recruited 16 participants (11f/5m, 0 other/non-binary, mean age= 24.94 years, SD = 2.19) with varying levels of experience in fabrication—ranging from complete novices to advanced users (see Table 3).
ID | Age | Experience | Group | Background/Description |
P1 | 26 | novice | G1 | Law background; no fabrication experience. |
P2 | 27 | advanced | G2 | CS background; experienced in laser cutting and 3D printing. |
P3 | 23 | basic | G3 | CS background; took hardware courses (Human Robot Interaction). |
P4 | 24 | basic | G4 | CS background; took hardware courses (Human Robot Interaction). |
P5 | 28 | novice | G1 | Information Science background; no fabrication experience. |
P6 | 27 | novice | G2 | Technology and product design background; no fabrication experience. |
P7 | 23 | novice | G3 | CS background; no fabrication experience. |
P8 | 23 | advanced | G4 | Architect; skilled in CAD/CAM with 6 years of experience. |
P9 | 28 | advanced | G1 | Fabrication researcher; CS expertise. |
P10 | 26 | advanced | G2 | MBA background; experience in pottery, crochet, and wood cutting. |
P11 | 27 | basic | G3 | Digital artist; trained in CAD (C4D, Blender) and laser cutting. |
P12 | 26 | advanced | G4 | Digital artist; proficient in laser cutting, 3D printing, CNC, and ceramic printing. |
P13 | 24 | advanced | G1 | Architect; completed independent fabrication projects. |
P14 | 20 | basic | G2 | Maker space worker; basic knowledge of processes. |
P15 | 24 | advanced | G3 | 6 years of design experience. |
P16 | 23 | basic | G4 | Maker space worker; basic knowledge of processes. |
Tasks. Participants were assigned three tasks: (Task 1) brainstorm and discuss methods to fabricate models they were asked to prepare/select before the workshop; (Task 2) then they studied alternative fabrication workflows using provided objects (shown in Figure 14) fabricated using different workflows, and discussed challenges and considerations when switching workflows; and (Task 3) used CAMeleon to collaboratively design and fabricate objects.
Procedure. The workshop took place in two sessions: one plenary session with all participants, which took 105 minutes, and follow-up fabrication sessions with the individual teams of 4. The plenary workshop began with a pre-workshop demographics survey taking 10min. Participants were randomly assigned to 4 teams of 4, we ensured each group had a mixture of participants with different levels of expertise. Task 1 took place in small groups and took 20min. Between Task 1 and 2 we gave a brief presentation explaining what a full fabrication workflow looks like, at the example of 3D printing. Task 2 was conducted within the same teams took and 20min, followed by Task 3 which took 30min. The workshop concluded with about 15min of short presentations of the different teams in which they shared their final workflows, considerations and discussed the role of CAMeleon. We scheduled a follow-up session with each team for the actual fabrication. Participants were free to take breaks and communicate with other teams.
Data collection. The goal was to understand the role and usefulness of CAMeleon during the design process. Researchers took notes during the different tasks to capture participants’ decision-making processes, challenges, and comparison of workflows. Participants were given whiteboards and paper to make notes, which have been captured by researchers, as well as continuous video running during the tasks.
During the follow-up fabrication session, we engaged in reflective semi-structured qualitative interviews with participants to gather additional insights about their experiences, challenges, and perceptions of CAMeleon’s support in completing their workflows.
10.1. Results
All teams successfully designed and fabricated their intended artifacts in the workshop as shown in Figure 15. Based on the notes, observations, and interviews during the workshop we report the following findings:
10.1.1. Supporting Users with Varying Skill Levels and Backgrounds
We observed that level of expertise played a big role in how users engaged with the presented tasks and CAMeleon. Here we break that down along the same dimensions of expertise we highlighted in the demographics.
Novices and Entry-Level Users. Beginners often participated using rather simple ”volume models.” These straightforward models allowed novices to focus on following CAMeleon’s guidance step-by-step without overthinking each part of the fabrication process. Observations showed this simplicity encouraged gradual exploration and experimentation with different fabrication methods. We found that CAMeleon flattened the learning curve by offering comparisons of workflows and materials and providing realistic previews. P1 (novice) found the tool accessible and helpful: ”I think I would definitely use it if there is such a need, …I think it is quite clear… you put the file in, and then drag it to get a preview, and you can roughly know what the result is, and there are labels below, and you can roughly know what materials are, which is quite interesting.” P6 (novice) appreciated that the software provides detailed step-by-step processes: ”The software providing the model gives a step-by-step detailed process, and the ability to compare different models is quite helpful.” Initially, when choosing a workflow, P6 did not think about specifics like material preferences. Still, by interacting with the model, they were guided to consider features like depth, width, and structure. We found that participants learned about unknown fabrication workflows as they explored, as P6 mentioned CAMeleon:”When I worked with the model, it guided me to think about the specific features I was looking for. This helped me plan better…By comparing different models—observing details such as depth, width, distinct layers, and overall structure—it helped me decide which design was easier and more convenient to build.”
Before using the software, group discussions often showed ”mutual influence.” If an advanced user proposes to use 3D printing, the inexperienced members will usually be attracted by this ”expert idea” and will directly accept the solution. This situation is mainly because Novice and Basic users do not have enough knowledge of other processes and cannot independently evaluate, which leads them to rely on the opinions of advanced users. However, when CAMeleon was introduced, the behavior of participants changed. They all explored the software process individually, and then the users returned to the team with the fabrication methods they had selected or explored in the software for group discussion. The discussion at this stage was more efficient and diverse because each participant had certain background information and personal insights, and the discussion was no longer limited to the single idea of the expert but extended to a higher level of ideas around the pros and cons of different processes.
This also indicated that: as participants engaged with CAMeleon, their initial ambiguity about fabrication workflows dissipated. For instance, participants collectively explored the trade-offs between workflows during group discussions, debating options like laser cutting versus 3D printing for specific tasks. This collaborative process enabled them to articulate the steps involved, compare methods, and weigh the pros and cons of each. To be more specific, reflecting on what the participants initially imagined the workflows looked like, they lacked many details. For example, during task 1, Group 4 specified that they would 3D scan an object and ”draw horizontal and vertical lines” on the model to use the laser cutter and 2D shaper to create the final artifact. There is a considerable gap between this imagined workflow and what happens in the fabrication workflow. Detailing these nuances in an interface enables these participants to understand the trade-offs of each workflow better and mitigate some of the ”in-situ” challenges that come up with less thoughtful designs.
CAMeleon’s interactive guidance function helps users get familiar with the fabrication process and increase their confidence. However, the fabrication process still requires practical experience, and guidance alone cannot bypass all difficulties. For example, P1 is a law student who quickly learned the steps of human-centered CAM by reading the interactive guide when using CAMeleon. This made P1 feel that they had mastered the process. However, P1 repeatedly encountered difficulties when replicating the fabrication workflow. In this case, P1 returned to the software to view the interactive guide again and re-understood every detail in the steps. Through repeated comparison and practice, P1 gradually realized that the ignored details were not dispensable and directly affected the operation’s success.
Users with Basic Fabrication Experience. For participants with basic fabrication knowledge, CAMeleon bridged the gap between digital design and fabrication. P11 highlighted CAMeleon’s ability to expand creative possibilities: ”I definitely think it can help. The biggest inspiration for me lies in the materials and fabrication aspect. For example, 3D modeling can be done in tools like AutoCAD, C4D, or Blender, but when it comes to bringing something to life, without this software, I might just to using 3D printing with some default material to handle it. However, with this software, it has expanded my thinking and made me consider more possibilities.” P11 also mentioned ”[CAMeleon] let them make appropriate comparisons before fabrication.” P4, who has experience in robotic hardware design, described its potential in prototyping and reducing trial-and-error: “When we designed soccer robots, we made a rough prototype to confirm feasibility before refining materials and details. If software could help confirm the entire workflow at the start, it would be more efficient.”
A noteworthy example was P11’s use of CAMeleon to innovate process combinations. Initially, P11 used laser cutting to create an internal support structure and applied clay to shape it. Later, inspired by the soft texture of needle felting workflow by reintroducing the model to CAMeleon, and following the interactive guidance to learning steps, P11 added a needle-felted layer on top of the clay structure. By using CAMeleon to iterate and fine-tune these workflows, P11 completed a project with rich layers and unique textures. This highlights how CAMeleon supports not only traditional workflows but also innovative explorations that transcend material and process limitations.
Advanced Users. Advanced participants had experience with CAD tools like Rhino or Grasshopper. For them, CAMeleon acted as a complementary tool to explore fabrication possibilities. P8, an architecture professional with six years of experience, described CAMeleon as: ’For me, it’s more like a dictionary tool to see what’s possible, and then I can use other software to further refine it.’ P8 also proposed two hybrid usage modes for advanced users: 1) As an ”Interactive Dictionary” CAMeleon helps users explore workflows, compare options, and visualize results. This enables users to identify suitable workflows early and transition to their preferred design tools (e.g., Rhino, Fusion) to build the workflow and detailed adjustments. 2) For Workflow Comparison and File Preparation CAMeleon streamlines early steps, such as generating machine-ready files (e.g., SVG, STL), which users can then refine in CAD tools to meet project-specific design or fabrication needs. This avoids manual setup and repetitive adjustments. P12, a digital craft professional, highlighted its utility for installations and decision making: ”I think it can definitely support me. When it comes to comparing workflows, it’s actually quite useful. It might be even more beneficial for designers, as they often need to compare different materials. For other cases, like creating [art] installations, this kind of comparison would also be helpful.”
The exploratory potential of CAMeleon is further exemplified by P8’s group work during the workshop. P8, an architect, originally prepared a model of what was called a “candelabra” (see Figure 16a). At that stage, P8 did not have a clear idea of how to fabricate the model or a definitive final design scheme. By importing the model into CAMeleon—despite the tool primarily targeting volumetric models and the initial design not fully meeting those criteria—and browsing different fabrication workflows to understand the requirements and limitations, P8 began an iterative exploration process.
Based on the insights from CAMeleon regarding the model’s needs, P8 refined the design in successive steps, exploring various workflows. This process led to successfully processing the model in CAMeleon (see Figure 16b) and exporting a lasercut-ready SVG. P8 then imported the SVG into Rhino (see Figure 16c) for further refinement, iteratively carving out a central space to accommodate a candle (see Figure 16d). Through this process, CAMeleon enabled P8 and the team not only to engage in exploration in fabrication, but also to iteratively refine the design—using feedback to inform choices and leveraging CAMeleon to visualize and confirm those decisions. The final result is shown in Figure 15c.
10.1.2. Promoting Exploration in Fabrication
P11 reflected on how her initial fabrication approach was constrained by familiarity, leaning heavily toward 3D printing: ”…If I were to fabricate my own model, I might not know what materials to use. Initially, I tended to lean toward 3D printing because I was relatively more familiar with that manufacturing process. I didn’t know much about other fabrication methods… But after attending the workshop, I realized it’s possible to use other materials, like cotton thread, paper, or bamboo… to incorporate into the model. This was something I hadn’t considered before.” P4 described CAMeleon as a valuable reference of existing workflows: “It broadens your thinking and provides instructions… I learned new methods.” Similarly, P8 emphasized the tool’s role in listing fabrication workflows:“It lists multiple possible fabrication solutions, helping me explore workflows I’m not familiar with. At the very least, it gives me an idea of what those workflows might looks like.”
In the initial brainstorming period, participants were not just thinking about fabrication process but also about the final artifact and its materiality. For example, on one of the whiteboards, Group 2 had initially envisioned creating a dog or cat. Rather than relying on plastic/wood as the primary materials for these fabrication workflows that may be easier, they felt drawn to softer materials that dictated the workflows that they ended up using(needle felting) and CAMeleon suggested having foam and wool to create a less rigid body. The material selection played a significant role in the design process, they indicated a preference for workflows that matched their resulting vision of how the artifact would feel. CAMeleon visualizes this for them and enables them to explore which material fits their vision best.
10.1.3. Combining Computational Fabrication and Craft
Participants observed a gap between traditional craftsmanship and digital fabrication, which they think CAMeleon could bridge. P11 highlighted the divide:”…Everyone’s [fabrication] mindset is too digitalized… There’s either no one exploring, or there’s this clear divide between traditional [fabrication] and digital. I can feel it traditional methods are very traditional, with their own logic that’s deeply rooted in tradition, while digital methods are completely different…However, digital objects often become repetitive and lose the unique touch of traditional materials…There’s no exploration of something in between.” Reflecting on a hybrid model combining a copper coating , bamboo which made by CAMeleon, P11 remarked:: ”That process was very novel to me…It expands my mindset and lets me integrate traditional crafts with modern technology,” P1 noted “I think there is a bit of a gap [between craftsmanship and digital fabrication]” From P1’s perspective, one major benefit of human-centered-CAM workflows is that “it’s like the machine helps me create a framework, and I just need to add material on top.” This not only saves time and materials but also increases creative freedom: “There are certain structures I definitely wouldn’t be able to make if I were just doing by hand.”
During Group 2’s discussion, while brainstorming how to create a “fluffy dog,” the group collaboratively sketched a voluminous design, aiming to emphasize the artifact’s soft, furry texture. Although this artifact inherently aligns with traditional handcraft techniques, such as felting or crochet, the group, including advanced maker P10 (who has extensive experience in pottery, crochet, and wood cutting), expressed a preference for integrating ”digital fabrication tools”—specifically, using an electric clipper to shape the structure. This preference underscores how even advanced makers, despite their strong backgrounds in manual techniques, recognize the limitations of purely traditional approaches when faced with specific challenges like precision shaping or structural consistency.
10.1.4. Interface and Enhancements
Participants liked the simple and straightforward design of CAMeleon. P9 noted:The software quickly takes one object and shows it in different ways…It lets me compare different manufacturing workflows very clearly. Similarly, P3 highlighted the interface: ”…Each option has detailed instructions, which gives users a clear expectation of what they can achieve.” P1 explains that ”Overall, I think the software is quite straightforward” and emphasizes that “most of the interactive guidance is really helpful for a beginner like me.”
However, as this is research software, there were plenty of feature requests to improve CAMeleon:
Integration with CAD Tools P9 noted the lack of integration with design tools as a limitation: “It is a standalone page and not integrated into design tools, which makes it less convenient. It would be better if I could directly modify the model.” P12 suggested that integrating CAMeleon with modeling software could allow users to adjust models during the design process, rather than repeatedly revisiting designs after encountering manufacturing issues: ”If it could integrate with other design software, like modeling tools… it could allow users to adjust the model during the design process to some extent, rather than having to repeatedly modify the model after completing the design due to manufacturing-related issues.” P8 suggested integration with tools like Grasshopper and Rhino to enhance compatibility with established workflows: “It would be much more convenient if the webpage could directly connect with these commonly used software tools.”
Improving Performance In the workshop, some users were dissatisfied with the running speed of CAMeleon, which affected their experience to a certain extent. Modules in CAMeleon currently are there to support a large variety of design workflows, and are not yet optimized for performance. When loading or manipulating larger models, this sometimes caused latency problems.
Other suggestions P4 proposed incorporating large language models (LLMs) to enable natural language interaction, particularly for less experienced users: “Getting faster, more intuitive feedback through natural language could make software easier to use.”. P4 suggested a feature to recommend alternative modules when specific steps fail or are infeasible: “If a workflow has several modules and one of them doesn’t work, it would be helpful if the software could recommend new methods or replacements to adjust the workflow quickly.”. P1 pointed out that even if CAMeleon can generate machine instructions (e.g., SVG or CSV), some beginners might still struggle: “…may still not be able to use the machine. Like how to adjust the settings, or preparing the document and then putting it into an app… I don’t really understand those details.”. P11 suggested adding more examples of printing successes and failures, saying that it ”If you could include more details, like examples of printing failures and successes, it would make things clearer for users without much experience.” P12 highlighted the value of showing practical data like “budget or weight,” especially for larger objects,“One suggestion I made earlier was to include more practical data, like budget or weight. Sometimes, when objects get larger, different materials can make them quite heavy.”
During the workshop we find that, advanced participants often approached CAMeleon with a predetermined mindset regarding fabrication workflows for different parts during the model design stage. For instance, some users had clearly planned, “This part should be 3D printed, and that part should be laser cut.” Currently, CAMeleon treats the entire model as a single object, even when it contains multiple components, which means users cannot assign different fabrication workflows to individual parts.
Additionally, the software does not allow for differentiation within a single model when elements of varying scales exist. An example is P15, who brought a sand table architectural model consisting of floor slabs, beams, walls, tables, and chairs. Parameters optimized for fabricating the walls at 10 cm in height could not simultaneously achieve sufficient detail for the chairs at less than 3 cm. Without the ability to separate workflows for different parts of the model, P15 found it impossible to balance fidelity and efficiency. Faced with these limitations, P15 simplified their model by removing intricate details like tables and chairs, retaining only the primary structures such as floors and walls. This allowed them to experiment with alternative fabrication workflows that better suited the simplified model. Expert functionality, such as supporting assemblies or segmenting geometry, could address these challenges in the future.
10.2. Discussion
Overall, our evaluation showed that CAMeleon played a leading role in participant’s design processes. It served as a platform for exploration, comparison, and education around different fabrication workflows.
10.2.1. Bridging the Gap Between Guidance and Practical Application
We found that CAMeleon and the Human-Centered CAM workflows enables users without knowledge about specific workflows to compare different options and even fabricate using workflows they may initially have been unaware of. However, as P1’s case revealed: theoretical knowledge alone does not fully prepare users for the nuances and unexpected variables that arise during actual fabrication. This highlights the importance of tacit knowledge in fabrication workflows—knowledge that extends beyond replicating steps and often defies easy articulation, as ’we can know more than we can tell’ (Polanyi, 1967). Future iterations of CAMeleon could provide real-time, contextual support, such as AR or voice-based guidance, to bridge this gap and help users navigate the dynamic and emergent nature of craft workflows.
10.2.2. Bridging Traditional Craftsmanship and Digital Fabrication
CAMeleon enables users to create works that retain the unique tactile qualities of handcraft while introducing innovative design possibilities. The introduction of Human-Centered CAM crucially extended this beyond what a machine can do, but explicitly embraces human effort in the export format. By offering these at the same level of fidelity, the workshop shows CAMeleon allowed designers to focus on personalized exploration. Even advanced users with expertise in fabrication embraced virtual comparisons, illustrating that hybrid workflows benefit both beginners and experienced makers. Moreover, it showed CAMeleon empowers rather than replaces traditional skills, fostering creativity and innovation such as remixing existing workflows (Zoran and Buechley, 2013).
10.2.3. Enhancing Integration in the Design-to-Fabrication Workflow
Advanced users of CAMeleon revisit the platform to review guidance before returning to their CAD tools to adjust designs. This pattern suggests a need for stronger support in the feedback loop between design and fabrication. Participants pointed this out as a clear opportunity for progress, albeit outside of the scope of this paper.
We see such tighter integration between CAMeleon and design environments as a fruitful step forwards, such integration could expose constraints of fabrication workflows in the design process and maybe nudge designers towards alternative fabrication workflows as they make decisions on their design. Generally, crossing the divide between CAD and CAM more naturally may be a powerful direction for the development of future tools.
10.2.4. Rethinking Workflow Choices Through Collective Exploration
Our observations show that workflow decisions are often influenced by experienced participants. By democratizing access to workflows, CAMeleon changed this pattern. This gave inexperienced users confidence to return to group discussions with their own insights, making the conversations more diverse and productive. The impact of CAMeleon goes beyond just individual exploration. It changed how participants viewed their roles in decision-making, allowing them to confidently challenge the status quo. This shifted the focus of discussions in the design process to consider a wider range of perspectives.
These findings suggest a potential for improving interdisciplinary collaboration. In many hardware design projects, each member is often working within the boundaries of their own expertise. Tools like CAMeleon could encourage experts from different fields to step outside of their specific expertise, fostering shared insights and more integrated problem-solving. This approach has the potential to create a collaborative environment where decisions are enriched by contributions from all perspectives, leading to more innovative and well-rounded outcomes.
11. CONCLUSION AND FUTURE WORK
We have presented CAMeleon a tool to explore and compare workflows for computational fabrication. We introduce the idea of human-centered CAM workflows as synergistic workflows between classical CAM and manual craft processes. In a workshop evaluation, we demonstrated CAMeleon’s utility and found that designing and exploring on a workflow level profoundly impacts the design process, and that it benefits designers at different levels of expertise. The modular architecture of CAMeleon supported our efforts to extend it with workflows from craftspeople and research projects evidences the extensibility of CAMeleon and helped identify future opportunities for technical implementation.
We see CAMeleon as a small step towards a future where fabrication users share and explore workflows and processes, instead of overly specific models for fabrication. This is similar to how recipe books for cooking serve as a starting point for inspiration instead of an endpoint to replicate the exact same dish.
Moving forward, we plan to advance the extensibility of CAMeleon by creating a visual programming environment where craftspeople without coding experience can remix new workflows and design their own. We furthermore plan to distribute CAMeleon in local fabrication and craft communities to study its longer-term usage and adoption.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.References
- (1)
- Goo (2010) 2010. Science, policy, and stakeholders in water management: an integrated approach to river basin management. Earthscan, London; Washington, DC.
- Alderighi et al. (2018) Thomas Alderighi, Luigi Malomo, Daniela Giorgi, Nico Pietroni, Bernd Bickel, and Paolo Cignoni. 2018. Metamolds: computational design of silicone molds. ACM Transactions on Graphics 37, 4 (Aug. 2018), 1–13. https://doi.org/10.1145/3197517.3201381
- Baudisch et al. (2017) Patrick Baudisch, Stefanie Mueller, and others. 2017. Personal fabrication. Foundations and Trends® in Human–Computer Interaction 10, 3–4 (2017), 165–293. Publisher: Now Publishers, Inc..
- Baudisch et al. (2023) Patrick Baudisch, Stefanie Mueller, Thijs Roumen, Pedro Lopes, Robert Kovacs, Jotaro Shigeyama, Shohei Katakura, Muhammad Abdullah, Conrad Lempert, Martin Taraz, and Lukas Rambold. 2023. Pushing Fabrication Research past the Makers. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–5. https://doi.org/10.1145/3544549.3583922
- Baudisch et al. (2019) Patrick Baudisch, Arthur Silber, Yannis Kommana, Milan Gruner, Ludwig Wall, Kevin Reuss, Lukas Heilman, Robert Kovacs, Daniel Rechlitz, and Thijs Roumen. 2019. Kyub: A 3D Editor for Modeling Sturdy Laser-Cut Objects. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–12. https://doi.org/10.1145/3290605.3300796
- Bourgault et al. (2023) Samuelle Bourgault, Pilar Wiley, Avi Farber, and Jennifer Jacobs. 2023. CoilCAM: Enabling Parametric Design for Clay 3D Printing Through an Action-Oriented Toolpath Programming System. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–16. https://doi.org/10.1145/3544548.3580745
- Derhamy et al. (2018) Hasan Derhamy, Mattias Andersson, Jens Eliasson, and Jerker Delsing. 2018. Workflow management for edge driven manufacturing systems. In 2018 IEEE Industrial Cyber-Physical Systems (ICPS). IEEE, St. Petersburg, 774–779. https://doi.org/10.1109/ICPHYS.2018.8390805
- Deshpande et al. (2021) Himani Deshpande, Haruki Takahashi, and Jeeeun Kim. 2021. EscapeLoom: Fabricating New Affordances for Hand Weaving. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–13. https://doi.org/10.1145/3411764.3445600
- Dumas et al. (2014) Jérémie Dumas, Jean Hergel, and Sylvain Lefebvre. 2014. Bridging the gap: automated steady scaffoldings for 3D printing. ACM Transactions on Graphics 33, 4 (July 2014), 1–10. https://doi.org/10.1145/2601097.2601153
- Fossdal et al. (2023) Frikk H Fossdal, Vinh Nguyen, Rogardt Heldal, Corie L. Cobb, and Nadya Peek. 2023. Vespidae: A Programming Framework for Developing Digital Fabrication Workflows. In Proceedings of the 2023 ACM Designing Interactive Systems Conference. ACM, Pittsburgh PA USA, 2034–2049. https://doi.org/10.1145/3563657.3596106
- Gao et al. (2023) Peizhong Gao, Tanhao Gao, Yanbin Yang, Zhenyuan Liu, Jianyu Shi, and Jin Li. 2023. Bamboo Agents: Exploring the Potentiality of Digital Craft by Decoding and Recoding Process. In Proceedings of the Seventeenth International Conference on Tangible, Embedded, and Embodied Interaction. ACM, Warsaw Poland, 1–13. https://doi.org/10.1145/3569009.3572746
- Hartikainen et al. (2024) Heidi Hartikainen, Leena Ventä-Olkkonen, Marta Cortés Orduña, Ivan Sanchez Milara, Marjukka Käsmä, and Leena Kuure. 2024. Making The Future School: An Analysis of Teens’ Collaborative Digital Fabrication Project. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 91:1–91:37. https://doi.org/10.1145/3637368
- Higgins et al. (2022) Erin Higgins, William Berkley Easley, Karen L. Gordes, Amy Hurst, and Foad Hamidi. 2022. Creating 3D Printed Assistive Technology Through Design Shortcuts: Leveraging Digital Fabrication Services to Incorporate 3D Printing into the Physical Therapy Classroom: Leveraging Digital Fabrication Services to Incorporate 3D Printing into the Physical Therapy Classroom. In Proceedings of the 24th International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’22). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3517428.3544816
- Hurst and Tobias (2011) Amy Hurst and Jasmine Tobias. 2011. Empowering individuals with do-it-yourself assistive technology. In The proceedings of the 13th international ACM SIGACCESS conference on Computers and accessibility. 11–18.
- Iarussi et al. (2015) Emmanuel Iarussi, Wilmot Li, and Adrien Bousseau. 2015. WrapIt: computer-assisted crafting of wire wrapped jewelry. ACM Transactions on Graphics 34, 6 (Nov. 2015), 1–8. https://doi.org/10.1145/2816795.2818118
- Ingold (2010) Tim Ingold. 2010. The textility of making. Cambridge journal of economics 34, 1 (2010), 91–102.
- McCrae et al. (2014a) James McCrae, Nobuyuki Umetani, and Karan Singh. 2014a. FlatFitFab: interactive modeling with planar sections. In Proceedings of the 27th annual ACM symposium on User interface software and technology. ACM, Honolulu Hawaii USA, 13–22. https://doi.org/10.1145/2642918.2647388
- McCrae et al. (2014b) James McCrae, Nobuyuki Umetani, and Karan Singh. 2014b. FlatFitFab: interactive modeling with planar sections. In Proceedings of the 27th annual ACM symposium on User interface software and technology. ACM, New York, NY, USA, 13–22.
- Mori and Igarashi (2007) Yuki Mori and Takeo Igarashi. 2007. Plushie: an interactive design system for plush toys. ACM Transactions on Graphics 26, 3 (July 2007), 45. https://doi.org/10.1145/1276377.1276433
- Nakashima et al. (2018) Kazutaka Nakashima, Thomas Auzinger, Emmanuel Iarussi, Ran Zhang, Takeo Igarashi, and Bernd Bickel. 2018. CoreCavity: interactive shell decomposition for fabrication with two-piece rigid molds. ACM Transactions on Graphics 37, 4 (Aug. 2018), 1–13. https://doi.org/10.1145/3197517.3201341
- Parry-Hill et al. (2017) Jeremiah Parry-Hill, Patrick C. Shih, Jennifer Mankoff, and Daniel Ashbrook. 2017. Understanding Volunteer AT Fabricators: Opportunities and Challenges in DIY-AT for Others in e-NABLE. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (CHI ’17). Association for Computing Machinery, New York, NY, USA, 6184–6194. https://doi.org/10.1145/3025453.3026045 event-place: Denver, Colorado, USA.
- Polanyi (1967) M. Polanyi. 1967. The Tacit Dimension. Anchor Books. https://books.google.com/books?id=jwLXAAAAMAAJ
- Rai et al. (2009) Sudhendu Rai, Charles B. Duke, Vaughn Lowe, Cyndi Quan-Trotter, and Thomas Scheermesser. 2009. LDP Lean Document Production—O.R.-Enhanced Productivity Improvements for the Printing Industry. Interfaces 39, 1 (Feb. 2009), 69–90. https://doi.org/10.1287/inte.1080.0413
- Roumen et al. (2022) Thijs Roumen, Ingo Apel, Thomas Kern, Martin Taraz, Ritesh Sharma, Ole Schlueter, Jeffrey Johnson, Dominik Meier, Conrad Lempert, and Patrick Baudisch. 2022. Structure-Preserving Editing of Plates and Volumes for Laser Cutting. In Proceedings of the 7th Annual ACM Symposium on Computational Fabrication. ACM, Seattle WA USA, 1–12. https://doi.org/10.1145/3559400.3561996
- Schiff and Bates (2010) Gordon D. Schiff and David W. Bates. 2010. Can Electronic Clinical Documentation Help Prevent Diagnostic Errors? New England Journal of Medicine 362, 12 (March 2010), 1066–1069. https://doi.org/10.1056/NEJMp0911734
- Schmidt and Umetani (2014) Ryan Schmidt and Nobuyuki Umetani. 2014. Branching support structures for 3D printing. In ACM SIGGRAPH 2014 Studio. ACM, Vancouver Canada, 1–1. https://doi.org/10.1145/2619195.2656293
- Schwartzburg and Pauly (2013) Yuliy Schwartzburg and Mark Pauly. 2013. Fabrication‐aware Design with Intersecting Planar Pieces. Computer Graphics Forum 32, 2pt3 (May 2013), 317–326. https://doi.org/10.1111/cgf.12051
- Sennett (2008) Richard Sennett. 2008. The craftsman. Yale University Press.
- Stava et al. (2012) Ondrej Stava, Juraj Vanek, Bedrich Benes, Nathan Carr, and Radomír Měch. 2012. Stress relief: improving structural strength of 3D printable objects. ACM Transactions on Graphics 31, 4 (Aug. 2012), 1–11. https://doi.org/10.1145/2185520.2185544
- Subbaraman and Peek (2022) Blair Subbaraman and Nadya Peek. 2022. p5.fab: Direct Control of Digital Fabrication Machines from a Creative Coding Environment. In Designing Interactive Systems Conference. ACM, Virtual Event Australia, 1148–1161. https://doi.org/10.1145/3532106.3533496
- Tibeme et al. (2018) Bernis Tibeme, Hossain Shahriar, and Chi Zhang. 2018. Process Mining Algorithms for Clinical Workflow Analysis. In SoutheastCon 2018. IEEE, St. Petersburg, FL, 1–6. https://doi.org/10.1109/SECON.2018.8479176
- Tojo et al. (2024) Kenji Tojo, Ariel Shamir, Bernd Bickel, and Nobuyuki Umetani. 2024. Fabricable 3D Wire Art. In Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24. ACM, Denver CO USA, 1–11. https://doi.org/10.1145/3641519.3657453
- Tran O’Leary et al. (2023) Jasper Tran O’Leary, Gabrielle Benabdallah, and Nadya Peek. 2023. Imprimer: Computational Notebooks for CNC Milling. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–15. https://doi.org/10.1145/3544548.3581334
- Tran O’Leary et al. (2024) Jasper Tran O’Leary, Thrisha Ramesh, Octi Zhang, and Nadya Peek. 2024. Tandem: Reproducible Digital Fabrication Workflows as Multimodal Programs. In Proceedings of the CHI Conference on Human Factors in Computing Systems. ACM, Honolulu HI USA, 1–16. https://doi.org/10.1145/3613904.3642751
- Tseng and Resnick (2014) Tiffany Tseng and Mitchel Resnick. 2014. Product versus process: representing and appropriating DIY projects online. In Proceedings of the 2014 conference on Designing interactive systems. ACM, Vancouver BC Canada, 425–428. https://doi.org/10.1145/2598510.2598540
- Tseng and Tsai (2015) Tiffany Tseng and Geoff Tsai. 2015. Process Products: Capturing Design Iteration with Digital Fabrication. In Proceedings of the Ninth International Conference on Tangible, Embedded, and Embodied Interaction. ACM, Stanford California USA, 631–636. https://doi.org/10.1145/2677199.2687891
- Valkeneers et al. (2019) Tom Valkeneers, Danny Leen, Daniel Ashbrook, and Raf Ramakers. 2019. StackMold: Rapid Prototyping of Functional Multi-Material Objects with Selective Levels of Surface Details. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology. ACM, New Orleans LA USA, 687–699. https://doi.org/10.1145/3332165.3347915
- Wall et al. (2021) Ludwig Wilhelm Wall, Alec Jacobson, Daniel Vogel, and Oliver Schneider. 2021. Scrappy: Using Scrap Material as Infill to Make Fabrication More Sustainable. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–12. https://doi.org/10.1145/3411764.3445187
- Zhao et al. (2018) Jian Zhao, Michael Glueck, Petra Isenberg, Fanny Chevalier, and Azam Khan. 2018. Supporting Handoff in Asynchronous Collaborative Sensemaking Using Knowledge-Transfer Graphs. IEEE Transactions on Visualization and Computer Graphics 24, 1 (Jan. 2018), 340–350. https://doi.org/10.1109/TVCG.2017.2745279
- Zoran and Buechley (2013) Amit Zoran and Leah Buechley. 2013. Hybrid reassemblage: an exploration of craft, digital fabrication and artifact uniqueness. Leonardo 46, 1 (2013), 4–10.
Appendix A Workflows Sources
Here is the list of workflows sources with links to their instruction pages.
Appendix B Step-by-Step Needle Felting Human-centered CAM Workflow